### v2

# Collocation methods for nonlinear differential equations on low-rank manifolds

Abstract
We introduce new methods for integrating nonlinear differential equations on low-rank manifolds. The methods are based on oblique projectors onto the tangent spaces of low-rank manifolds with an interpolation property that collocates the differential equation on the manifold. These projections yield evolution equations that enable low-rank time integration of any vectorfield that can be evaluated entry-wise. The proposed methods do not require the vectorfield to have a low-rank structure, thus overcoming significant limitations of dynamical low-rank approximation based on orthogonal projection. To construct the oblique projectors, we devise a sparse tensor sampling algorithm based on the discrete empirical interpolation method (DEIM) that parameterizes tensor train manifolds and their tangent spaces with cross interpolation. Using these projectors we propose two time integration schemes on low-rank tensor train manifolds. The first is based on a time-dependent tensor cross interpolation representation of the solution and the second is a direct generalization of the well-known orthogonal projector-splitting integrator to a oblique projectors. We demonstrate the proposed methods with applications to several tensor differential equations arising from the discretization of partial differential equations.
keywords:
low-rank approximation , time-dependent tensors , tensor differential equations , tensor cross approximation , tensor train format
journal: Arxiv
1 Introduction
Consider the initial value problem
$\frac{\partial u(\bm{x},t)}{\partial t}=\mathcal{G}({u},\bm{x},t),\qquad{u}(\bm{x},0)={u}_{0}(\bm{x}),$ (1)
governing the time evolution of a quantity of interest $u:\Omega\times[0,T]\to\mathbb{R}$, where $\Omega$ is a subset of $\mathbb{R}^{d}$ ($d\gg 1$) and $\mathcal{G}$ is a nonlinear operator that may depend on $\bm{x}$ and $t$. Equations of the form (1) are found in many areas of physical sciences, engineering, and mathematics. For example, in applications of kinetic theory such as the Fokker–Planck equation [41] and the Boltzmann equation [4], in optimal mass transport [19], and as finite-dimensional approximations of functional differential equations [53, 52]. Discretizing (1) with a method of lines yields the tensor differential equation
$\frac{dX(t)}{dt}=G(X(t),t),\qquad X(0)=X_{0},$ (2)
where $X(t):[0,T]\to\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ is the time-dependent solution tensor and $G:\mathbb{R}^{n_{1}\times\cdots\times n_{d}}\times[0,T]\to\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ is a discrete form of the operator $\mathcal{G}$. At any time $t$, the solution tensor $X(t)$ has $\mathcal{O}(n^{d})$ degrees of freedom that make its computation and storage prohibitively expensive, even for small $d$.
Several algorithms based on tensor networks have recently been proposed to reduce the number of degrees of freedom in the solution tensor $X(t)$ and integrate (2) at a reasonable computational cost. A tensor network is a factorization of a high-dimensional tensor, such as $X(t)$, into a network of low-dimensional tensors with significantly fewer degrees of freedom. The number of degrees of freedom in a network depends on the chosen tensor format, e.g., tensor train (TT) [38], Tucker [34, 13], Hierarchical Tucker [22, 33, 23] or canonical polyadic (CP) [29], and the tensor rank. For instance, tensors in the TT format with rank $r$ can be parameterized with $\mathcal{O}(dnr^{2})$ degrees of freedom, a significant reduction from $\mathcal{O}(n^{d})$ when the rank $r$ is sufficiently small. The set of all tensors in a chosen format with fixed rank forms a smooth manifold on which the solution to (2) can be integrated.
Two classes of algorithms for integrating (2) on smooth tensor manifolds are step-truncation methods [42] and dynamical low-rank methods [10]. Step-truncation methods allow the solution rank to naturally increase in a controlled manner during a time step before truncating back to the desired rank. Dynamical low-rank methods integrate the solution on a fixed-rank manifold by projecting $G(X,t)$ onto a tangent space of the manifold at each time $t$. Both methods aim to compute the best approximate solution to (2) on a fixed-rank tensor manifold at each time $t$ and are consistent with each other as the temporal step-size approaches zero (see, e.g., [9, Section 3.3]). These methods have been utilized for several applications including uncertainty quantification [2, 45], plasma physics [54, 17], numerical approximation of functional differential equations [53, 44] and machine learning [48, 46], and substantial research efforts have recently been made to improve their accuracy, efficiency, and robustness. These efforts have resulted in several innovations including rank-adaptive integrators [9, 5, 42], implicit low-rank methods [43, 36, 50], conservative low-rank methods [25, 3, 18, 16] and coordinate-adaptive low-rank methods [11, 12].
Despite these recent advancements, existing low-rank time integration schemes still have significant limitations in their applicability. Most notably, they require the tensor-valued map $G$ defining the differential equation (2) to have low-rank structure complementary to that of the solution. Such low-rank structure is key to increasing the solution rank in a controlled manner for step-truncation methods or efficiently computing the orthogonal projection of $G(X,t)$ onto the tangent space for dynamical low-rank methods. In either case, the low-rank structure of $G$ is crucial for obtaining practical time integration schemes with computational cost and storage requirements comparable to the storage cost of the chosen tensor format. However, many instances of $G$ that arise from discretizing (1) lack low-rank structure. For example, when $G$ includes a polynomial nonlinearity computing a low-rank representation of $G(X,t)$ or its orthogonal projection onto the tangent space is expensive due to the non-optimal rank that results from multiplying low-rank tensors. The situation is worse for other common nonlinearities, such as exponential or fractional, as there are currently no reliable algorithms for performing these nonlinear arithmetic operations with low-rank tensors. In such cases, (2) may admit an approximate low-rank solution. However, existing step-truncation and dynamical low-rank methods cannot efficiently compute it.
In this paper, we introduce a new class of dynamical low-rank methods that can efficiently integrate the solution to (2) on a low-rank tensor manifold even when $G$ does not have low-rank structure. Our proposed methods rely on a new class of oblique projectors onto low-rank tangent spaces with a cross interpolation property. In the context of dynamical low-rank approximation, these projectors collocate (2) on a tensor manifold and yield equations of motion on the manifold that are efficient to integrate whenever it is possible to evaluate $G$ entry-wise. The oblique projectors are defined by sets of multi-indices that identify tensor fibers along which the projector interpolates. To select these multi-indices, we introduce a new algorithm, based on the discrete empirical interpolation method (DEIM), to efficiently compute indices that parameterize tensor manifolds with cross interpolation [37]. Using these projectors we propose two low-rank time integration schemes. The first integrates subtensors of the solution defining a tensor cross interpolant and then constructs the low-rank solution later in time with tensor cross interpolation. The second we obtain by applying a splitting scheme to the oblique tangent space projector. Splitting schemes for orthogonal tangent space projectors were introduced in [31] for matrices and were subsequently generalized to TTs [32], Tucker tensors [30] and tree tensor networks [7]. Our method directly generalizes these orthogonal projector-splitting schemes to oblique projectors in the TT format. Related time integration schemes based on oblique projections onto the low-rank manifold (instead of its tangent space) were recently proposed for computing low-rank approximations to matrix differential equations [15, 35] and tensor differential equations [21, 20].
The rest of this paper is organized as follows. In Section 2 we briefly recall the TT format and orthogonalization of tensors in the TT format. In Section 3 we recall the orthogonal projector onto the TT tangent space and introduce new oblique projectors onto the tangent space. Then we describe a special class of oblique projectors with a cross interpolation property. In Section 4 we present a new index selection algorithm, referred to as TT-cross-DEIM, for constructing oblique projectors onto the TT tangent space. We show that indices obtained with the TT-cross-DEIM algorithm define well-conditioned oblique tangent space projectors and parameterize TT manifolds with cross interpolation. In Section 5 we introduce new dynamical low-rank time integration schemes for (2) using oblique tangent space projectors. In Section 6 we demonstrate the proposed dynamical low-rank methods and compare the results with existing time integration methods on low-rank tensor manifolds. The main findings are summarized in Section 7.
2 Tensor train format
To introduce notation and make our presentation self-contained we begin with a brief review of the tensor train (TT) format and orthogonal TT representations. For more details on the TT format we refer the reader to [38]. The $k$th unfolding of a tensor $Y\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ is the matrix $\bm{Y}^{\langle k\rangle}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times(n_{k+1}\cdots n_{d})}$ obtained by reshaping the entries of $Y$, where the row and column indices form a colexicographical ordering. The TT-rank of $Y$ is defined as the vector $\bm{r}=(1,r_{1},\ldots,r_{d},1)$ where $r_{k}$ is the rank of the unfolding matrix $\bm{Y}^{\langle k\rangle}$. Any tensor $Y$ with TT-rank $\bm{r}$ can be represented in the TT format as
$Y(i_{1},i_{2},\ldots,i_{d})=C_{1}(i_{1})C_{2}(i_{2})\cdots C_{d}(i_{d}),$ (3)
where each $C_{k}$ is a $r_{k-1}\times n_{k}\times r_{k}$ tensor referred to as a TT-core and $C_{k}(i_{k})$ is a $r_{k-1}\times r_{k}$ matrix for a fixed index $i_{k}$. Each TT-core has a left unfolding matrix $\bm{C}_{k}^{\langle l\rangle}\in\mathbb{R}^{r_{k-1}n_{k}\times r_{k}}$ and a right unfolding matrix $\bm{C}_{k}^{\langle r\rangle}\in\mathbb{R}^{r_{k-1}\times n_{k}r_{k}}$ obtained by reshaping the elements of $C_{k}$
$\bm{C}_{k}^{\langle l\rangle}(\alpha_{k-1}i_{k},\alpha_{k})=\bm{C}_{k}^{\langle r\rangle}(\alpha_{k-1},i_{k}\alpha_{k})=C_{k}(\alpha_{k-1},i_{k},\alpha_{k}),\qquad k=1,2,\ldots,d.$
The left and right unfoldings of each TT-core is full rank whenever $Y$ has TT-rank $\bm{r}$.
            (a)                                                                  (b)
Figure 1: Tensor network diagrams of a $d=6$ dimensional tensor in the TT format. (a) No orthogonalization with the TT-core partial products $C_{\leq 2}$ and $C_{>2}$ indicated. (b) Orthogonalized TT (6) with $k=2$ and left orthogonal TT-cores $U_{\leq 2}$ and right orthogonal TT-cores $V_{>2}$ indicated.
To simplify notation of tensors in the TT format we often omit indices so that (3) is replaced by $Y=C_{1}C_{2}\cdots C_{d}$. To obtain even more compact representations of tensors in the TT format, we define partial products of TT-cores $C_{\leq k}\in\mathbb{R}^{n_{1}\times\cdots\times n_{k}\times r_{k}}$ and $C_{>k}\in\mathbb{R}^{r_{k}\times n_{k+1}\times\cdots\times n_{d}}$ with entries
$\begin{split}C_{\leq k}(i_{1},\ldots,i_{k},:)&=C_{1}(i_{1})\cdots C_{k}(i_{k},:),\\ C_{>k}(:,i_{k+1},\ldots,i_{d})&=C_{k+1}(:,i_{k+1})\cdots C_{d}(i_{d}),\end{split}$ (4)
so that $Y=C_{\leq k}C_{>k}$. We also define certain unfolding matrices of these partial product tensors
$\begin{split}\bm{C}_{\leq k}(i_{1}\cdots i_{k},:)&=C_{\leq k}(i_{1},\ldots,i_{k},:),\\ \bm{C}_{>k}(i_{k+1}\cdots i_{d},:)&=C_{>k}(:,i_{k+1},\ldots,i_{d}),\end{split}$ (5)
where $\bm{C}_{\leq k}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times r_{k}}$ and $\bm{C}_{>k}\in\mathbb{R}^{(n_{k+1}\cdots n_{d})\times r_{k}}$ to write the $k$th unfolding matrix of $Y$ as $\bm{Y}^{\langle k\rangle}=\bm{C}_{\leq k}\bm{C}_{>k}^{\text{T}}$.
2.1 Orthogonalization of tensor trains
Orthogonal TT representations are fundamental for executing many operations in the TT format. We will use them in this paper to obtain well-conditioned projectors onto the tangent spaces of TT manifolds. Hereafter, we recall an algorithm for orthogonalizing TTs by recursively applying QR-decomposition to TT-core unfoldings. Begin by taking a QR-decomposition of the left unfolding of $C_{1}$
$\bm{C}_{1}^{\langle l\rangle}=\bm{U}_{1}^{\langle l\rangle}\bm{R}_{1},$
to obtain the matrix $\bm{R}_{1}\in\mathbb{R}^{r_{1}\times r_{1}}$ and the new TT-core $U_{1}\in\mathbb{R}^{r_{0}\times n_{1}\times r_{1}}$ defined by its left unfolding. The new TT-core is called left-orthogonal because it satisfies
$\left[\bm{U}_{1}^{\langle l\rangle}\right]^{\text{T}}\bm{U}_{1}^{\langle l\rangle}=\bm{I}_{r_{1}},$
where $\bm{I}_{r_{1}}$ denotes the $r_{1}\times r_{1}$ identity matrix. Next define a new second core $\widehat{C}_{2}(i_{2})=\bm{R}_{1}C_{2}(i_{2})$ to obtain the TT representation $Y=U_{1}\widehat{C}_{2}C_{>2}$ of the tensor (3). Then take a QR-decomposition of the left unfolding of the second core
$\widehat{\bm{C}}_{2}^{\langle l\rangle}=\bm{U}_{2}^{\langle l\rangle}\bm{R}_{2},$
to obtain $\bm{R}_{2}\in\mathbb{R}^{r_{2}\times r_{2}}$ and the left-orthogonal TT-core $U_{2}\in\mathbb{R}^{r_{1}\times n_{2}\times r_{2}}$. Define a new third core $\widehat{C}_{3}(i_{3})=\bm{R}_{2}C_{3}(i_{3})$ to write $Y=U_{\leq 2}\widehat{C}_{3}C_{>3}$ where now the first two cores are left orthogonal. Proceeding recursively in this way we obtain the TT representation
$Y=U_{\leq k}\bm{R}_{k}C_{>k},$
where $U_{j}$ is left-orthogonal for $j=1,2,\ldots,k$ and $\bm{R}_{k}\in\mathbb{R}^{r_{k}\times r_{k}}$. Similar to orthogonalizing cores from left to right, we can also orthogonalize cores from right to left by recursively performing QR-decompositions on right unfoldings (see [38, Section 3]) to obtain
$Y=U_{\leq k}\bm{R}_{k}\bm{R}_{k+1}V_{>k},$
where $\bm{R}_{k+1}\in\mathbb{R}^{r_{k}\times r_{k}}$ and the $V_{j}$ are right-orthogonal, i.e.,
$\bm{V}_{j}^{\langle r\rangle}\left[\bm{V}_{j}^{\langle r\rangle}\right]^{\text{T}}=\bm{I}_{r_{j-1}},\qquad j=k+1,\ldots,d.$
Letting $\bm{S}_{k}=\bm{R}_{k}\bm{R}_{k+1}$ we obtain the orthogonalized TT representation
$Y=U_{\leq k}\bm{S}_{k}V_{>k},$ (6)
and utilizing the unfolding matrices of partial products (5) we also have a decomposition of the $k$th unfolding
$\bm{Y}^{\langle k\rangle}=\bm{U}_{\leq k}\bm{S}_{k}\bm{V}_{>k}^{\text{T}}.$ (7)
It follows from the left orthogonality of $U_{j}$ and right orthogonality of $V_{j}$ that the columns of $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ are orthonormal, i.e., $\bm{U}_{\leq k}^{\text{T}}\bm{U}_{\leq k}=\bm{V}_{>k}^{\text{T}}\bm{V}_{>k}=\bm{I}_{r_{k}}$. The decomposition (7) resembles a SVD however $\bm{S}_{k}$ is not necessarily diagonal. If the TT-rank of $Y$ is $\bm{r}$ then $\bm{S}_{k}$ is invertible. What is important for the projectors defined in the subsequent section is that the columns of $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ form orthonormal bases for the range and co-range of $\bm{Y}^{\langle k\rangle}$ respectively.
3 Projections onto the TT tangent space
It is well-known that the collection of all rank-$\bm{r}$ TTs
$\mathcal{M}_{\bm{r}}=\{Y\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}\mid\text{TT-rank}(Y)=\bm{r}\},$ (8)
is a smooth embedded submanifold of $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ [27]. Hence for any tensor $Y\in\mathcal{M}_{\bm{r}}$ we can define the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ as the collection of (equivalence classes of) velocities of smooth curves on $\mathcal{M}_{\bm{r}}$ passing through $Y$
$T_{Y}\mathcal{M}_{\bm{r}}=\left\{\gamma^{\prime}(s)|_{s=0}\mid\gamma(s)\in C^{1}((-\delta,\delta),\mathcal{M}_{\bm{r}}),\ \gamma(0)=Y\right\}.$ (9)
The tangent space is a vector subspace of $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ that linearizes the manifold $\mathcal{M}_{\bm{r}}$ around $Y$. Given a TT representation (3) of $Y$, any element of the tangent space has the form
$\delta Y=\delta C_{1}C_{2}\cdots C_{d}+C_{1}\delta C_{2}C_{3}\cdots C_{d}+\cdots+C_{1}\cdots C_{d-1}\delta C_{d},$ (10)
where $\delta C_{k}\in\mathbb{R}^{r_{k-1}\times n_{k}\times r_{k}}$ are first order variations of the TT-cores.
3.1 The orthogonal tangent space projector
The orthogonal projector $\widehat{P}_{Y}:\mathbb{R}^{n_{1}\times\cdots\times n_{d}}\to T_{Y}\mathcal{M}_{\bm{r}}$ onto the tangent space considered in [32] determines the best approximation of a given tensor $Z\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ in the tangent space relative to the Frobenius norm. Such orthogonal projector can be constructed from the orthogonal projectors
$\begin{split}\widehat{\bm{P}}_{\leq k}=\bm{U}_{\leq k}\bm{U}_{\leq k}^{\text{T}},\qquad\widehat{\bm{P}}_{>k}=\bm{V}_{>k}\bm{V}_{>k}^{\text{T}},\qquad k=1,2,\ldots,d-1,\end{split}$ (11)
onto the range and co-range of $\bm{Y}^{\langle k\rangle}$ (see (7)). These projectors act on the matrix space $\mathbb{R}^{(n_{1}\cdots n_{k})\times(n_{k+1}\cdots n_{d})}$. To construct the orthogonal tangent space projector it is convenient to define projectors corresponding to (11) that act on the tensor space $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ by
$\begin{split}\widehat{P}_{\leq k}Z=\text{Ten}_{k}\left[\widehat{\bm{P}}_{\leq k}\bm{Z}^{\langle k\rangle}\right],\qquad\widehat{P}_{>k}Z=\text{Ten}_{k}\left[\bm{Z}^{\langle k\rangle}\widehat{\bm{P}}_{>k}\right],\end{split}$ (12)
where $\text{Ten}_{k}$ denotes the tensorization operator that is the inverse of the $k$th unfolding, i.e., $\text{Ten}_{k}\left(\bm{Z}^{\langle k\rangle}\right)=Z$. The projectors $\widehat{P}_{\leq j},\widehat{P}_{>k}$ commute whenever $j\leq k$ and can be used to construct the orthogonal projector onto the tangent space [32, Corollary 3.2]
$\widehat{P}_{Y}=\sum_{k=1}^{d-1}\widehat{P}_{\leq k-1}\widehat{P}_{>k}-\widehat{P}_{\leq k}\widehat{P}_{>k}+\widehat{P}_{\leq d-1},$ (13)
where we set $\widehat{P}_{\leq 0}=1$.
3.2 Oblique tangent space projectors
The computational cost of dynamical low-rank approximation of (2) with orthogonal tangent space projections can scale as $\mathcal{O}(n^{d})$ when $G$ lacks low-rank structure. In this case, the orthogonal dynamical low-rank method is impractical as its computational cost is comparable to solving (2) without low-rank compression. To enable efficient dynamical low-rank approximation in such cases, we introduce oblique projections onto the TT tangent space that can be computed in only $\mathcal{O}(dnr^{3})$ operations for many applications where orthogonal projections require $\mathcal{O}(n^{d})$. We construct such oblique projectors by replacing the orthogonal projectors defined in (11) with oblique projectors onto the same spaces
$\begin{split}{\bm{P}}_{\leq k}=\bm{U}_{\leq k}\left(\bm{X}^{\text{T}}_{\leq k}\bm{U}_{\leq k}\right)^{-1}\bm{X}^{\text{T}}_{\leq k},\qquad{\bm{P}}_{>k}=\bm{X}_{>k}\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)^{-\text{T}}\bm{V}_{>k}^{\text{T}},\end{split}$ (14)
defined for any matrices $\bm{X}_{\leq k}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times r_{k}}$ and $\bm{X}_{>k}\in\mathbb{R}^{(n_{k+1}\cdots n_{d})\times r_{k}}$ such that $\left(\bm{X}^{\text{T}}_{\leq k}\bm{U}_{\leq k}\right)$ and $\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)$ are invertible. Just as before, it is convenient to define oblique projectors corresponding to (14) that act on the tensor space $\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ by
$\begin{split}{P}_{\leq k}Z=\text{Ten}_{k}\left[\bm{P}_{\leq k}\bm{Z}^{\langle k\rangle}\right],\qquad{P}_{>k}Z=\text{Ten}_{k}\left[\bm{Z}^{\langle k\rangle}\bm{P}_{>k}\right].\end{split}$ (15)
With these projectors we can construct an oblique tangent space projector with the same form as the orthogonal tangent space projector (13).
Figure 2: A sketch of the low-rank tensor manifold $\mathcal{M}_{\bm{r}}$ and its tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ at the point $Y\in\mathcal{M}_{\bm{r}}$. Also depicted is a tensor $Z\in\mathbb{R}^{n_{1}\times\cdots n_{d}}$ and its orthogonal projection $\widehat{P}_{Y}Z$ and oblique projection $P_{Y}Z$ onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. The orthogonal projection is the best approximation of $Z$ on the tangent space with respect to the Frobenius norm but is impractical compute when $G$ lacks rank structure. The oblique projection is a quasi-optimal approximation of $Z$ on the tangent space that is efficient to compute for any $G$ that can be evaluated entry-wise.
Proposition 3.1
Let $Y\in\mathcal{M}_{\bm{r}}$ with orthogonal decompositions of its unfolding matrices given in (7) and suppose $\bm{X}_{\leq k},\bm{X}_{>k}$ define oblique projectors (14). Then the map
${P}_{Y}=\sum_{k=1}^{d-1}{P}_{\leq k-1}{P}_{>k}-{P}_{\leq k}{P}_{>k}+{P}_{\leq d-1},$ (16)
with $P_{\leq 0}=1$, defined for any tensor $Z\in\mathbb{R}^{n_{1}\times\cdots n_{d}}$, is an oblique projector onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$.
Proof: First we show that the image of $P_{Y}$ is contained in $T_{Y}\mathcal{M}_{\bm{r}}$. Note that $T_{Y}\mathcal{M}_{\bm{r}}$ is a linear space and thus it is sufficient to show that the image of each term in (16) belongs to $T_{Y}\mathcal{M}_{\bm{r}}$. To do so we utilize [32, Corollary 3.2] which shows that $P_{\leq j},P_{>k}$ commute whenever $j\leq k$ and
$P_{\leq k-1}P_{>k}Z=\text{Ten}_{k}\left\{\left[\bm{I}_{n_{k}}\otimes\bm{P}_{\leq k-1}\right]\bm{Z}^{\langle k\rangle}\bm{P}_{>k}\right\}.$ (17)
Note that the result is stated for orthogonal projectors $\widehat{P}_{\leq k},\widehat{P}_{>k}$ however the proof does not require orthogonality of the projectors and thus holds for oblique projectors ${P}_{\leq k},{P}_{>k}$. Inserting the oblique projectors (14) into (17) we obtain
$\begin{split}P_{\leq k-1}P_{>k}Z&=\text{Ten}_{k}\left\{\left[\bm{I}_{n_{k}}\otimes\bm{U}_{\leq k-1}\left(\bm{X}^{\text{T}}_{\leq k-1}\bm{U}_{\leq k-1}\right)^{-1}\bm{X}^{\text{T}}_{\leq k-1}\right]\bm{Z}^{\langle k\rangle}\left[\bm{X}_{>k}\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)^{-T}\bm{V}_{>k}^{\text{T}}\right]\right\}\\ &=\text{Ten}_{k}\left\{\left[\bm{I}_{n_{k}}\otimes\bm{U}_{\leq k-1}\right]\left[\bm{I}_{n_{k}}\otimes\left(\bm{X}^{\text{T}}_{\leq k-1}\bm{U}_{\leq k-1}\right)^{-1}\bm{X}^{\text{T}}_{\leq k-1}\bm{Z}^{\langle k\rangle}\bm{X}_{>k}\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)^{-T}\right]\bm{V}_{>k}^{\text{T}}\right\}\\ &=\text{Ten}_{k}\left\{\left[\bm{I}_{n_{k}}\otimes\bm{U}_{\leq k-1}\right]\delta\bm{C}_{k}^{\langle l\rangle}\bm{V}_{>k}^{\text{T}}\right\}\end{split}$ (18)
where we used the mixed product property of the Kronecker product to obtain the second equality and defined $\delta C_{k}$ as the $r_{k-1}\times n_{k}\times r_{k}$ TT-core with left unfolding
$\delta\bm{C}_{k}^{\langle l\rangle}=\bm{I}_{n_{k}}\otimes\left(\bm{X}^{\text{T}}_{\leq k-1}\bm{U}_{\leq k-1}\right)^{-1}\bm{X}^{\text{T}}_{\leq k-1}\bm{Z}^{\langle k\rangle}\bm{X}_{>k}\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)^{-T}.$ (19)
Using the identity
$\bm{C}_{\leq k}=\left(\bm{I}_{n_{k}}\otimes\bm{C}_{\leq k-1}\right)\bm{C}_{k}^{\langle l\rangle},$ (20)
and (5) we can write (18) in TT format
$P_{\leq k-1}P_{>k}Z=U_{\leq k-1}\delta C_{k}V_{>k},$ (21)
which has the form (10) and thus belongs to the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. For the $P_{\leq k}P_{>k}$ terms we have
$\begin{split}P_{\leq k}P_{>k}Z&=\text{Ten}_{k}\left\{\bm{P}_{\leq k}\bm{Z}^{\langle k\rangle}\bm{P}_{>k}\right\}\\ &=\text{Ten}_{k}\left\{\bm{U}_{\leq k}\left[\left(\bm{X}^{\text{T}}_{\leq k}\bm{U}_{\leq k}\right)^{-1}\bm{X}^{\text{T}}_{\leq k}\bm{Z}^{\langle k\rangle}\bm{X}_{>k}\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)^{-T}\right]\bm{V}_{>k}^{\text{T}}\right\},\end{split}$ (22)
which we write in the TT format
$P_{\leq k}P_{>k}Z=U_{\leq k}\delta\bm{S}_{k}V_{>k},$ (23)
where
$\delta\bm{S}_{k}=\left(\bm{X}^{\text{T}}_{\leq k}\bm{U}_{\leq k}\right)^{-1}\bm{X}^{\text{T}}_{\leq k}\bm{Z}^{\langle k\rangle}\bm{X}_{>k}\left(\bm{X}_{>k}^{\text{T}}\bm{V}_{>k}\right)^{-T},$ (24)
belongs to $\mathbb{R}^{r_{k}\times r_{k}}$. Absorbing $\delta\bm{S}_{k}$ into its left neighboring core $U_{k}$ we rewrite (23) as
$P_{\leq k}P_{>k}Z=U_{\leq k-1}\delta\overline{C}_{k}V_{>k},$ (25)
where $\delta\overline{C}_{k}(i_{k})=U_{k}(i_{k})\delta\bm{S}_{k}$, which has the form (10) and therefore belongs to the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$.
It remains to show that $P_{Y}$ is idempotent, which we verify by checking that $P_{\leq k}$ and $P_{>k}$ are idempotent for all $k=1,\ldots,d-1$. For $P_{\leq k}$ we have
$\begin{split}{P}_{\leq k}^{2}Z&=P_{\leq k}\text{Ten}_{k}\left[\bm{P}_{\leq k}\bm{Z}^{\langle k\rangle}\right]\\ &=\text{Ten}_{k}\left[\bm{P}_{\leq k}^{2}\bm{Z}^{\langle k\rangle}\right]\\ &=\text{Ten}_{k}\left[\bm{P}_{\leq k}\bm{Z}^{\langle k\rangle}\right]\\ &={P}_{\leq k}Z,\end{split}$ (26)
for all $k=1,\ldots,d-1$. The idempotence of $P_{>k}$ for $k=1,\ldots,d-1$ is also easy to verify directly. $\square$
We remark that the orthogonality of the bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ is not necessary for the construction of oblique tangent space projectors (16). However, imposing such orthogonality is advantageous for the numerical stability and accuracy of the projectors. Next we introduce a special class of oblique tangent space projectors with a cross interpolation property.
3.3 Interpolatory tangent space projectors
Let us consider oblique tangent space projectors (16) with $\bm{X}_{\leq k}$ and $\bm{X}_{>k}$ defined as $r_{k}$ columns of the identity matrix with appropriate dimension
$\begin{split}\bm{X}_{\leq k}=\bm{I}_{n_{1}\cdots n_{k}}\left(:,I^{\leq k}\right),\qquad\bm{X}_{>k}=\bm{I}_{n_{k+1}\cdots n_{d}}\left(:,I^{>k}\right).\end{split}$ (27)
Here, $I^{\leq k}$ contains $r_{k}$ indices of the form $i_{1}\cdots i_{k}$ and $I^{>k}$ contains $r_{k}$ indices of the form $i_{k+1}\cdots i_{d}$. The matrix $\bm{X}_{\leq k}$ can extract $r_{k}$ rows determined by the index sets $I^{\leq k}$ from a tensor $\bm{Z}^{\langle k\rangle}$ with matrix multiplication from the left $\bm{X}_{\leq k}^{\text{T}}\bm{Z}^{\langle k\rangle}=\bm{Z}^{\langle k\rangle}(I^{\leq k},:)$. Similarly, the matrix $\bm{X}_{>k}$ can extract $r_{k}$ columns from $\bm{Z}^{\langle k\rangle}$ with matrix multiplication from the right $\bm{Z}^{\langle k\rangle}\bm{X}_{>k}=\bm{Z}^{\langle k\rangle}(:,I^{>k}).$ Moreover, when $\bm{X}_{\leq k},\bm{X}_{>k}$ are defined as in (27) the oblique projectors in (14) become interpolatory projectors
$\begin{split}\left[\bm{P}_{\leq k}\bm{Z}^{\langle k\rangle}\right]\left(I^{\leq k},:\right)=\bm{Z}^{\langle k\rangle}\left(I^{\leq k},:\right),\qquad\left[\bm{Z}^{\langle k\rangle}\bm{P}_{>k}\right]\left(:,I^{>k}\right)=\bm{Z}^{\langle k\rangle}\left(:,I^{>k}\right).\end{split}$ (28)
Since each index $i_{1}\cdots i_{k}$ corresponds to a multi-index $(i_{1},\ldots,i_{k})$ and $i_{k+1}\cdots i_{d}$ corresponds to a multi-index $(i_{k+1},\ldots,i_{d})$, we identify the sets of $r_{k}$ indices $I^{\leq k}$ and $I^{>k}$ with sets of $r_{k}$ multi-indices $\mathcal{I}^{\leq k}$ and $\mathcal{I}^{>k}$. With such identification the matrix interpolation property (28) is equivalent to an interpolation property of the corresponding tensor operators (15)
$\begin{split}\left[P_{\leq k}Z\right]\left(\mathcal{I}^{\leq k},i_{k+1},\ldots,i_{d}\right)=Z\left(\mathcal{I}^{\leq k},i_{k+1},\ldots,i_{d}\right),\qquad\left[P_{>k}Z\right]\left(i_{1},\ldots,i_{k},\mathcal{I}^{>k}\right)=Z\left(i_{1},\ldots,i_{k},\mathcal{I}^{>k}\right).\end{split}$ (29)
We obtain interpolatory tangent space projectors (16) by considering multi-indices that satisfy the nested conditions
$\mathcal{I}^{\leq k}\subset\mathcal{I}^{\leq k-1}\times\{1,\ldots,n_{k}\},\qquad\mathcal{I}^{>k}\subset\{1,\ldots,n_{k}\}\times\mathcal{I}^{>k+1},\quad k=1,2,\ldots,d-1.$ (30)
To prove that (30) is sufficient for the oblique tangent space projector (16) to have a cross interpolation property we have the following Lemma.
Lemma 3.1
For any nested indices (30) defining oblique projectors (15) and $k=1,2,\ldots,d-1$, the projector $P_{\leq k-1}P_{>k}$ satisfies
$\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=\begin{cases}Z\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right),&j=k,\\ \left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right),&j>k,\\ \left[P_{\leq k-1}P_{>k-1}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right),&j<k.\end{cases}$ (31)
Proof: The case $j=k$ follows directly from the interpolation property (29). For $j>k$ the nesting condition (30) ensures that the first $k-1$ indices of each multi-index in $\mathcal{I}^{\leq j-1}$ is a multi-index in $\mathcal{I}^{\leq k-1}$ and that the first $k$ indices of each multi-index in $\mathcal{I}^{\leq j-1}$ is a multi-index in $\mathcal{I}^{\leq k}$. Therefore we can use the interpolation property (29) to obtain
$\left[P_{\leq k-1}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\ldots,i_{d}\right)=\left[P_{\leq k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\ldots,i_{d}\right)=Z\left(\mathcal{I}^{\leq j-1},i_{j},\ldots,i_{d}\right).$ (32)
Since $P_{\leq k}$ and $P_{\leq k-1}$ act only on the first $k$ dimensions of $Z$ and $P_{>k}$ acts only on dimensions $k+1,\ldots,d$ we can use the preceding equation to obtain
$\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=\left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right),$
establishing the case $j>k$. When $j<k$ the nested condition ensures that the $d-j$ indices of each multi-index in $\mathcal{I}^{>j}$ appear in multi-indices belonging to $\mathcal{I}^{>k}$ and $\mathcal{I}^{>k-1}$. Therefore from the interpolation property (29) we have
$\left[P_{>k}Z\right]\left(i_{1},\ldots,i_{j},\mathcal{I}^{>j}\right)=\left[P_{>k-1}Z\right]\left(i_{1},\ldots,i_{j},\mathcal{I}^{>j}\right)=Z\left(i_{1},\ldots,i_{j},\mathcal{I}^{>j}\right),$ (33)
from which we obtain the result for $j<k$. $\square$
Theorem 3.1
For any $Y\in\mathcal{M}_{\bm{r}}$ and $\{\mathcal{I}^{\leq j},\mathcal{I}^{>j}\}$ nested multi-indices defining interpolatory projectors (14) the oblique tangent space projector (16) has the cross interpolation property
$\left[P_{Y}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=Z\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right),\qquad j=1,2,\ldots,d.$ (34)
Proof: Rearrange the terms in (16) to write
$\begin{split}\left[P_{Y}Z\right]&=\sum_{k=1}^{d}P_{\leq k-1}P_{>k}Z-\sum_{k=1}^{d-1}P_{\leq k}P_{>k}Z,\end{split}$ (35)
with $P_{>d}=1$ and evaluate at the indices $\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)$
$\begin{split}\left[P_{Y}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)&=\sum_{k=1}^{d}\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)-\sum_{k=1}^{d-1}\left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right).\end{split}$ (36)
Applying Lemma 3.1 to each term in the first summation in (36) we obtain
$\begin{split}&\sum_{k=1}^{d}\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=\\ &\qquad\left(\sum_{k=1}^{j-1}\left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},:,\mathcal{I}^{>j}\right)\right)+Z\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)+\left(\sum_{l=j+1}^{d}\left[P_{\leq l-1}P_{>l-1}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)\right)\end{split}$ (37)
Re-indexing the final summation in (37) with $k=l-1$ and combining the result with the first summation in (37) yields
$\begin{split}&\sum_{k=1}^{d}\left[P_{\leq k-1}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)=Z\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right)+\sum_{k=1}^{d-1}\left[P_{\leq k}P_{>k}Z\right]\left(\mathcal{I}^{\leq j-1},i_{j},\mathcal{I}^{>j}\right).\end{split}$ (38)
Finally substituting (38) into (36) the two summations cancel and the proof is complete. $\square$
The cross interpolation property (34) of the tangent space projector resembles the cross interpolation property of TT-cross approximation [47] with nested indices. Next, we describe a novel index selection algorithm for constructing well-conditioned interpolatory projectors onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. Then we show that the resulting indices parameterize $Y$ with tensor cross interpolation.
4 Index selection for oblique projectors and cross interpolation
In the preceding section we constructed tangent space projectors (16) with the cross interpolation property (34) from oblique projectors (14) onto the bases $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$. We now devise an efficient algorithm based on the discrete empirical interpolation method (DEIM) for computing indices $\{I^{\leq k},I^{>k}\}$, or equivalently multi-indices $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$, that yield well-conditioned oblique projectors (14) defined by the matrices (27). The DEIM (recalled in Algorithm 1) is a greedy algorithm for computing indices that yield well-conditioned interpolatory projectors onto a given basis, e.g., $\bm{U}_{\leq k}$ or $\bm{V}_{>k}$. However we can not apply such algorithm directly to the matrices $\bm{U}_{\leq k},\bm{V}_{>k}$ as they have dimensions $(n_{1}\cdots n_{k})\times r_{k}$ and $(n_{k+1}\cdots n_{d})\times r_{k}$, respectively, which is too large to store in memory and process with the DEIM algorithm. To address the problem of memory, we do not store the matrices $\bm{U}_{\leq k}$ and $\bm{V}_{>k}$ directly but rather the left orthogonal cores $U_{j}$ and right orthogonal cores $V_{j}$ that can be used to construct these matrices
$\begin{split}\bm{U}_{\leq k}(i_{1}\cdots i_{k},:)&=U_{1}(i_{1})\cdots U_{k}(i_{k},:),\\ \bm{V}_{>k}(i_{k+1}\cdots i_{d},:)&=V_{k+1}(:,i_{k+1})\cdots V_{d}(i_{d}),\qquad k=1,2,\ldots,d-1.\end{split}$ (39)
To address the computational cost we propose an algorithm that only samples from a small subset of entries of the matrices $\bm{U}_{\leq k},\bm{V}_{>k}$. The key idea is to compute the indices $I^{\leq k}$ recursively for $k=1,2,\ldots,d-1$ by sampling from $\bm{U}_{\leq k}$ only indices corresponding to multi-indices $\mathcal{I}^{\leq k}$ that are nested (30). By considering only nested indices we reduce the number of possible indices $i_{1}\cdots i_{k}\in I^{\leq k}$ from $n_{1}\cdots n_{k}$ to $r_{k-1}n_{k}$. We use the same idea to construct the nested index sets $\mathcal{I}^{>k}$ sequentially for $k=d-1,d-2,\ldots,1$. Remarkably, it is always possible to obtain well-conditioned interpolatory projectors using this nested sampling approach. Moreover, since the indices obtained from this sampling approach are nested by construction, the resulting tangent space projector (16) has the cross interpolation property (34). Such cross interpolation property will be useful for the dynamical low-rank approximation schemes presented in Section 5. Hereafter we present the nested index selection algorithm in detail and show that the resulting indices yield well-conditioned interpolatory projectors.
Algorithm 1 DEIM index selection (adapted from [49])
1:$\bm{V}\in\mathbb{R}^{n\times r}$ with $n\geq r$
2:$\bm{l}$, a vector with $r$ distinct indices from $\{1,\ldots,n\}$
3:$\bm{v}=\bm{V}(:,1)$
[…]
5:for $j=2,3,\ldots,r$  do
6:     $\bm{v}=\bm{V}(:,j)$
7:     $\bm{c}=\bm{V}(l,1:j-1)^{-1}\bm{v}(\bm{l})$
8:     $\bm{r}=\bm{v}-\bm{V}(:,1:j-1)\bm{c}$
9:     $[~{},l_{j}]=\max(|\bm{r}|)$
10:     $\bm{l}=[\bm{l};l_{j}]$
11:end for
4.1 The TT-cross-DEIM algorithm
Let us denote by $\texttt{DEIM}(\cdot)$ a subroutine that takes a matrix of size $n\times r$ as input and outputs a collection of $r$ indices computed with the DEIM index selection Algorithm 1. To compute the indices $I^{\leq j}$, begin by applying the DEIM algorithm to the $n_{1}\times r_{1}$ matrix $\bm{U}_{\leq 1}$
$I^{\leq 1}=\texttt{DEIM}\left(\bm{U}_{\leq 1}\right).$ (40)
To obtain $I^{\leq 2}$, construct the $r_{1}n_{2}\times r_{2}$ matrix
$\begin{split}\widehat{\bm{U}}_{\leq 2}(\alpha_{1}i_{2},\alpha_{2})=\bm{U}_{\leq 2}\left(I_{\alpha_{1}}^{\leq 1}i_{2},\alpha_{2}\right),\end{split}$ (41)
which is the restriction of $\bm{U}_{\leq 2}$ to the indices $I^{\leq 1}$ with $I^{\leq 1}_{\alpha_{1}}$ denoting the $\alpha_{1}$st index in $I^{\leq 1}$. Then sample $r_{2}$ indices from this restricted matrix
$\bm{l}_{\leq 2}=\texttt{DEIM}\left(\widehat{\bm{U}}_{\leq 2}\right).$ (42)
Due to the construction of $\widehat{\bm{U}}_{\leq 2}$ in (41), each index $\bm{l}_{\leq 2}(\alpha_{2})$ corresponds to a multi-index $(\bm{p}_{\leq 1}(\alpha_{2}),\bm{p}_{2}(\alpha_{2}))$ where $\bm{p}_{\leq 1}(\alpha_{2})$ identifies an index in $\mathcal{I}^{\leq 1}$ and $\bm{p}_{2}(\alpha_{2})$ identifies an index in $\{1,\ldots,n_{2}\}$. Construct the multi-index set
$\mathcal{I}^{\leq 2}_{\alpha_{2}}=\left(\mathcal{I}^{\leq 1}_{\bm{p}_{\leq 1}(\alpha_{2})},\bm{p}_{2}(\alpha_{2})\right),\qquad\alpha_{2}=1,2,\ldots,{r_{2}},$ (43)
which corresponds to the set of indices $I^{\leq 2}$. The remaining sets of indices $I^{\leq j}$ are obtained inductively with a similar procedure. After computing ${I}^{\leq j-1}$, construct the $r_{j-1}n_{j}\times r_{j}$ matrix
$\begin{split}\widehat{\bm{U}}_{\leq j}(\alpha_{j-1}i_{j},\alpha_{j})=\bm{U}_{\leq j}\left(I^{\leq j-1}_{\alpha_{j-1}}i_{j},\alpha_{j}\right),\end{split}$ (44)
which is the restriction of $\bm{U}_{\leq j}$ to the indices $I^{\leq j-1}$. Then sample $r_{j}$ indices from the restricted matrix
$\bm{l}_{\leq j}=\texttt{DEIM}\left(\widehat{\bm{U}}_{\leq j}\right).$ (45)
Due to the construction of $\widehat{\bm{U}}_{\leq j}$ in (44) each index $\bm{l}_{\leq j}(\alpha_{j})$ corresponds to a multi-index $(\bm{p}_{\leq j-1}(\alpha_{j}),\bm{p}_{j}(\alpha_{j}))$ where $\bm{p}_{\leq j-1}(\alpha_{j})$ identifies a multi-index in $\mathcal{I}^{\leq j-1}$ and $\bm{p}_{j}(\alpha_{j})$ identifies an index in $\{1,\ldots,n_{j}\}$. Construct the set $\mathcal{I}^{\leq j}$ with multi-indices
$\mathcal{I}^{\leq j}_{\alpha_{j}}=\left(\mathcal{I}^{\leq j-1}_{\bm{p}_{\leq j-1}(\alpha_{j})},\bm{p}_{j}(\alpha_{j})\right),\qquad\alpha_{j}=1,2,\ldots,{r_{j}},$ (46)
which corresponds to $I^{\leq j}$. This procedure computes index sets $I^{\leq j}$ sequentially for $j=1,2,\ldots,d-1$ by applying the DEIM sampling algorithm to the restricted matrices $\widehat{\bm{U}}_{\leq j}$ of dimension $r_{j-1}n_{j}\times r_{j}$. In practice the restricted matrices $\widehat{\bm{U}}_{\leq j}$ are not obtained from $\bm{U}_{\leq j}$ as written in (44) but rather from the low-dimensional tensor cores $U_{j}$ that construct $\bm{U}_{\leq j}$ in (39). The computation of $I^{\leq j}$ from the tensor cores $U_{j}$ is depicted in Figure 3 with $d=4$.
Figure 3: Illustration of the TT-cross-DEIM left-to-right sweep that computes multi-indices $\mathcal{I}^{\leq j}$ sequentially for $j=1,2,\ldots,d-1$ with $d=4$.
We compute the index sets $I^{>j}$ in a similar manner. First obtain $I^{>d-1}$ by sampling the $n_{d}\times r_{d-1}$ matrix $\bm{V}_{>d}$
$I^{>d-1}=\texttt{DEIM}\left(\bm{V}_{>d-1}\right).$ (47)
Then construct index sets $I^{>j}$ inductively for $j=d-2,d-3,\ldots,1$ as follows. After computing $I^{>j+1}$, construct the $n_{j+1}r_{j+1}\times r_{j}$ matrix $\widehat{\bm{V}}_{>j}$
$\begin{split}\widehat{\bm{V}}_{>j}\left(i_{j+1}\alpha_{j+1},\alpha_{j}\right)=\bm{V}_{>j}\left(i_{j+1}I^{>j+1}_{\alpha_{j+1}},\alpha_{j}\right)\end{split}$ (48)
which is the restriction of $\bm{V}_{>j}$ to the indices $\mathcal{I}^{>j+1}$. Then sample $r_{j}$ indices from the restricted matrix
$\bm{l}_{>j}=\texttt{DEIM}\left(\widehat{\bm{V}}_{>j}\right).$ (49)
Due to the construction of $\widehat{\bm{V}}_{>j}$ in (48) each index $\bm{l}_{>j}(\alpha_{j})$ corresponds to a multi-index $(\bm{p}_{j+1}(\alpha_{j}),\bm{p}_{>{j+1}}(\alpha_{j}))$, where $\bm{p}_{j+1}(\alpha_{j})$ identifies an index in $\{1,\ldots,n_{j+1}\}$ and $\bm{p}_{>{j+1}}(\alpha_{j})$ identifies a multi-index in $\mathcal{I}^{>j+1}$. Construct the set $\mathcal{I}^{>j}$ with multi-indices
$\mathcal{I}^{>j}_{\alpha_{j}}=\left(\bm{p}_{j+1}(\alpha_{j}),\mathcal{I}^{>j+1}_{\bm{p}_{>j+1}(\alpha_{j})}\right),\qquad\alpha_{j}=1,2,\ldots,{r_{j}},$ (50)
which corresponds to $I^{>j}$. Just as in the computation of $I^{\leq j}$, the restricted matrices $\widehat{\bm{V}}_{>j}$ are not obtained from $\bm{V}_{>j}$ as written in (48) but rather from the low-dimensional tensor cores $V_{j}$ that construct $\bm{V}_{>j}$ in (39).
The entire algorithm is summarized in Algorithm 2 and a tensor network diagram of the left-to-right sweep for computing $I^{\leq j}$ from the TT-cores $U_{j}$ is shown in Figure 3. In Algorithm 2 we denote by ind2sub the Matlab function that reshapes linear indices to multi-indices. As mentioned above, the multi-index sets ${\mathcal{I}^{\leq k},\mathcal{I}^{>k}}$ obtained in (46) and (50) are nested (30) by construction. Thus the oblique tangent space projector (16) constructed from these indices is a cross interpolant (see Theorem 3.1), provided it is well-defined. It will be shown shortly that such projector is in fact well-defined.
Algorithm 2 TT-cross-DEIM index selection
1:
2:$\bm{U}_{\leq j},\bm{V}_{>j}$, $j=1,\ldots,d-1$, orthogonal bases for range and co-range of $\bm{Y}^{\langle j\rangle}$ as in (7)
3:
4:$\{\mathcal{I}^{\leq j},\mathcal{I}^{>j}\}$, nested multi-index sets defining oblique projectors (14) via (27)
5:$I^{\leq 1}=\texttt{DEIM}\left(\bm{U}_{\leq 1}\right)$triangleright left-to-right sweep: computing $I^{\leq j}$
6:for $j=2$ to $d-1$  do
7:     $\widehat{\bm{U}}_{\leq j}=\bm{U}_{\leq j}\left(I^{\leq j-1}:,:\right)$
8:     $\bm{l}_{\leq j}=\texttt{DEIM}(\widehat{\bm{U}}_{j})$
9:     $\bm{p}_{\leq j-1},\bm{p}_{j}=\texttt{ind2sub}([r_{j-1},n_{j}],\bm{l}_{\leq j})$
10:     $\mathcal{I}^{\leq j}_{\alpha_{j}}=\left(\mathcal{I}^{\leq j-1}_{\bm{p}_{\leq j-1}(\alpha_{j})},\bm{p}_{j}(\alpha_{j})\right),\qquad\alpha_{j}=1,2,\ldots,r_{j}$
11:end for
12:$I^{>d-1}=\texttt{DEIM}\left(\bm{V}_{>d-1}\right)$triangleright right-to-left sweep: computing $I^{>j}$
13:for $j=d-2$ to $1$  do
14:     $\widehat{\bm{V}}_{>j}=\bm{V}_{>j}\left(:I^{>j+1},:\right)$
15:     $\bm{l}_{>j}=\texttt{DEIM}\left(\widehat{\bm{V}}_{>j}\right)$
16:     $\bm{p}_{j+1},\bm{p}_{>j+1}=\texttt{ind2sub}([n_{j},r_{j}],\bm{l}_{>j})$
17:     $\mathcal{I}^{>j}_{\alpha_{j}}=\left(\bm{p}_{j+1}(\alpha_{j}),\mathcal{I}^{>j+1}_{\bm{p}_{>j+1}(\alpha_{j})}\right),\qquad\alpha_{j}=1,2,\ldots,r_{j}$
18:end for
Computational cost
To simplify the operation count of the proposed TT-cross-DEIM algorithm, we assume that $r_{k}=r$ and $n_{k}=n$ for all $k=1,2,\ldots,d$. The Algorithm requires access to all $d-1$ orthogonal representations (6) which can be computed in $\mathcal{O}(dnr^{3})$ operations [38]. With these orthogonal representations available, each index set $I^{\leq k}$ and $I^{>k}$ for $k=1,2,\ldots,d-1$ is computed by applying DEIM to a matrix of size $rn\times r$ each of which requires $\mathcal{O}(nr)$ operations. Therefore the total number of operations in the TT-cross-DEIM algorithm is $\mathcal{O}(dnr^{3})$.
Note that the operation count of TT-cross-DEIM is dominated by the computation of orthogonal TT representations and thus has the same complexity as many common TT algorithms, e.g., TT rounding.
4.2 Condition of the oblique projectors
For the oblique projectors (14) to be defined (and thus for the oblique tangent space projector (16) to be defined) the $r_{j}\times r_{j}$ matrices $\bm{M}_{j}=\bm{X}_{\leq j}^{\text{T}}\bm{U}_{\leq j}$ and $\bm{N}_{j}=\bm{X}_{>j}^{\text{T}}\bm{V}_{>j}$ must be invertible. For interpolatory projectors, i.e., when $\bm{X}_{\leq j}$ and $\bm{X}_{>j}$ are defined as in (27), these matrices are given entry-wise by
$\bm{M}_{j}(\alpha_{j},\beta_{j})=\bm{U}_{\leq j}\left(I^{\leq j}_{\alpha_{j}},\beta_{j}\right),\qquad\bm{N}_{j}(\alpha_{j},\beta_{j})=\bm{V}_{>j}\left(I^{>j}_{\alpha_{j}},\beta_{j}\right).$ (51)
The next result shows that the TT-cross-DEIM produces indices that yield invertible matrices (51) and therefore define oblique projectors (14) and (16).
Theorem 4.1
If $Y\in\mathcal{M}_{\bm{r}}$ and $I^{\leq j},I^{>j}$ are obtained with the TT-cross-DEIM then the $r_{j}\times r_{j}$ matrices (51) are invertible for all $j=1,2,\ldots,d-1$.
Proof: We prove the result for $\bm{M}_{j}$ by induction on $j$. The $r_{1}$ indices $I^{\leq 1}$ are obtained in (40) from $\bm{U}_{\leq 1}$ with the DEIM. Because $Y\in\mathcal{M}_{\bm{r}}$, the unfolding matrix $\bm{U}_{\leq 1}$ is full rank. Thus we have from [49, Lemma 3.1] that the $r_{1}\times r_{1}$ matrix $\bm{M}_{1}=\bm{U}_{\leq 1}\left(I^{\leq 1},:\right)$ is full rank, establishing the result for $\bm{M}_{j}$ when $j=1$. Now assume that $\bm{M}_{j-1}$ is full rank. Rewriting (44) using (4)-(5) we have
$\begin{split}\widehat{\bm{U}}_{j}(\alpha_{j-1}i_{j},\alpha_{j})&=\bm{U}_{\leq j-1}\left(I^{\leq j-1}_{\alpha_{j-1}},:\right)\bm{U}_{j}^{\langle r\rangle}(:,i_{j}\alpha_{j})\\ &=\bm{M}_{j-1}(\alpha_{j-1},:)\bm{U}_{j}^{\langle r\rangle}(:,i_{j}\alpha_{j}).\end{split}$ (52)
By assumption $\bm{M}_{j-1}$ is full rank and since $Y\in\mathcal{M}_{\bm{r}}$ both unfoldings of $\bm{U}_{j}$ are full rank. It follows that $\widehat{\bm{U}}_{j}$ is full rank. In (45) the $r_{j}$ indices $\bm{l}_{\leq j}$ are obtained from $\widehat{\bm{U}}_{j}$ with the DEIM and invoking [49, Lemma 3.1] we have that the $r_{j}\times r_{j}$ matrix $\widehat{\bm{U}}_{j}(\bm{l}_{\leq j},:)$ is full rank. The indices $\mathcal{I}^{\leq j}$ are obtained from $\bm{l}_{\leq j}$ in (46) so that
$\bm{U}_{\leq j}\left(I^{\leq j}_{\alpha_{j}},\beta_{j}\right)=\widehat{\bm{U}}_{\leq j}\left(\bm{l}_{\leq j}(\alpha_{j}),\beta_{j}\right),$ (53)
proving the result for $\bm{M}_{j}$. The statement for $\bm{N}_{j}$ is proven similary with induction on $j=d-1,d-2,\ldots,1$. $\square$
In addition to producing invertible matrices (51), the TT-cross-DEIM is a greedy algorithm that keeps the condition number of such matrices as small as possible subject to the constraint that the indices are nested. Indeed, the DEIM algorithm is employed to obtain $\bm{l}_{\leq j}$ in (45) and $\bm{l}_{>j}$ in (49) which greedily minimizes the condition number of $\widehat{\bm{U}}_{\leq j}(\bm{l}_{\leq j},:)$ and $\widehat{\bm{V}}_{>j}\left(\bm{l}_{>j},:\right)$. Given that $\widehat{\bm{U}}_{\leq j}$ is the restriction of $\bm{U}_{\leq j}$ to indices $I^{\leq j-1}$ (see (44)) and $\widehat{\bm{V}}_{>j}$ is the restriction of $\bm{V}_{>j}$ to indices $I^{>j+1}$ (see (48)), the TT-cross-DEIM does in fact select $\{I^{\leq j},I^{>j}\}$ to minimize the condition number of (51) subject to the constraint that the corresponding multi-indices are nested (30).
4.3 Tensor cross interpolation
We have shown above that the multi-indices obtained with the TT-cross-DEIM yield a well-conditioned interpolatory projector (16) onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. Incidentally, we can use the same multi-index sets to parameterize $Y$ with tensor cross interpolation. Recall that TT-cross approximation [14, 37, 47, 40] is a specific instance of the TT format (3) that generalizes matrix CUR decomposition to tensors. In this representation the TT-cores are defined by the entries of $Y$
$\tilde{Y}(i_{1},\ldots,i_{d})=\prod_{k=1}^{d-1}Y\left(\mathcal{I}^{\leq k-1},i_{k},\mathcal{I}^{>k}\right)\left[Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k}\right)\right]^{-1}Y\left(\mathcal{I}^{\leq d-1},i_{d}\right),$ (54)
where for convenience we set $\mathcal{I}^{\leq 0}=\emptyset$. It is well-known that the nested condition (30) is sufficient for the tensor cross approximation (54) to be a tensor cross interpolant [47]
$\tilde{Y}\left(\mathcal{I}^{\leq k-1},i_{k},\mathcal{I}^{>k}\right)=Y\left(\mathcal{I}^{\leq k-1},i_{k},\mathcal{I}^{>k}\right),\qquad k=1,2,\ldots,d.$ (55)
We now use the nested multi-index sets constructed by the TT-cross-DEIM to prove that any TT can be exactly represented as a TT-cross interpolant. This result follows as a Corollary of Theorem 4.1.
Corollary 4.1
Any $Y\in\mathcal{M}_{\bm{r}}$ can be exactly represented as a rank-$\bm{r}$ TT-cross interpolant with nested indices.
Proof: Let $\{I^{\leq j},I^{>j}\}$ be nested index sets obtained with the TT-cross-DEIM. Using (7) write the $r_{j}\times r_{j}$ matrix $Y\left(\mathcal{I}^{\leq j},\mathcal{I}^{>j}\right)$ as
$\begin{split}Y\left(\mathcal{I}^{\leq j},\mathcal{I}^{>j}\right)&=\bm{U}_{\leq j}\left(I^{\leq j},:\right)\bm{S}_{j}\bm{V}_{>j}\left(I^{>j},:\right)^{\text{T}}\\ &=\bm{M}_{j}\bm{S}_{j}\bm{N}_{j}^{\text{T}},\end{split}$ (56)
where we used (51) to obtain the second equality. We have shown in Theorem 4.1 that $\bm{M}_{j}$ and $\bm{N}_{j}$ are invertible and since $Y\in\mathcal{M}_{\bm{r}}$, the matrices $\bm{S}_{j}$ are also invertible for all $j=1,2,\ldots,d-1$. Therefore the matrices $Y\left(\mathcal{I}^{\leq j},\mathcal{I}^{>j}\right)$ are invertible for all $j=1,2,\ldots,d-1$ and hence by [14, Theorem 2] the nested multi-indices $\{\mathcal{I}^{\leq j},\mathcal{I}^{>k}\}$ provide an exact representation of $Y$ with TT-cross interpolation. $\square$
Several algorithms for computing tensor cross approximations (54) from black-box tensors based on the maximum volume principle have recently been developed [37, 14]. The purpose of our proposed TT-cross-DEIM algorithm is to obtain well-conditioned interpolatory tangent space projections for a tensor $Y\in\mathcal{M}_{\bm{r}}$, not for black-box tensor approximation. However, it was recently demonstrated in [20] that DEIM-based cross approximation algorithms can be applied iteratively to obtain tensor cross approximations from black-box tensors with comparable performance to the corresponding maximum volume algorithms.
5 Time integration on low-rank tensor manifolds
In this section we introduce dynamical low-rank schemes based on oblique tangent space projections for approximating the nonlinear tensor differential equation (2) on the TT manifold (8). The idea of dynamical low-rank methods is to integrate an approximate solution $Y(t)$ on $\mathcal{M}_{\bm{r}}$ by projecting $G(Y,t)$ defining the ODE (2) onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ at each time $t$. Conventionally, dynamical low-rank methods have employed orthogonal projectors (13) to obtain the best approximation of $G(Y,t)$ in the tangent space (see Figure 2). However, as we mentioned previously, the computational cost of orthogonal projection onto the tangent space can be $\mathcal{O}(n^{d})$ when $G$ lacks low-rank structure. We now propose new dynamical low-rank methods based on oblique tangent space projections that have cost scaling as $\mathcal{O}(dnr^{3})$ in for many applications where methods based on orthogonal projections cost $\mathcal{O}(n^{d})$. The proposed oblique dynamical low-rank methods are efficient whenever it is possible to evaluate $G(X,t)$ entry-wise and do not require $G$ to have any low-rank structure.
If the initial condition $X_{0}$ in (2) belongs to $\mathcal{M}_{\bm{r}}$, then the solution to the projected tensor differential equation
$\frac{dY(t)}{dt}=P_{Y(t)}G(Y(t),t),\qquad Y(0)=X_{0},$ (57)
is a dynamical low-rank approximation of the solution to (2) that remains on $\mathcal{M}_{\bm{r}}$ for all $t$. Hereafter we propose two time integration schemes for solving the dynamical low-rank equation (57) with interpolatory tangent space projectors (16). The first scheme, referred to as TT-cross time integration, integrates forward in time the entries of the solution tensor $Y(t)$ required to construct the TT-cross interpolant (54) at any time $t$. The second scheme is a direct generalization of the projector-splitting integrator for orthogonal dynamical low-rank approximation introduced in [31] for matrices and subsequently generalized to TTs [32], Tucker tensors [30] and tree tensor networks [7], to oblique tangent space projectors in the TT format.
5.1 TT-cross time integration
The time-dependent oblique tangent space projector in (57) is defined at each time $t$ by a set of time-dependent multi-indices $\{\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t)\}$. We assume that such multi-indices are nested (30) at each time $t$ so that the oblique projector has the cross interpolation property (34) at each $t$. The evolution equations defining the TT-cross integrator we obtain by directly evaluating (57) at the multi-indices $\{\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t)\}$ and utilizing the cross interpolation property (34)
$\frac{dY\left(\mathcal{I}^{\leq k-1}(t),:,\mathcal{I}^{>k}(t),t\right)}{dt}=G_{Y}\left(\mathcal{I}^{\leq k-1}(t),:,\mathcal{I}^{>k}(t),t\right),\vspace{.2cm}\qquad k=1,2,\ldots,d,$ (58)
where we defined the tensor $G_{Y}(t)=G\left(Y(t),t\right)$. Note that (58) is a set of evolution equations for a subset of entries of the approximate solution $Y(t)\in\mathcal{M}_{\bm{r}}$ that define a TT-cross approximation (54) at any given time $t$, hence the name TT-cross integrator. If $G$ arises from the spatial discretization of a PDE (1) involving differential operators then evaluating $G_{Y}\left(\mathcal{I}^{\leq k-1},:,\mathcal{I}^{>k}\right)$ requires entries of $Y$ at indices adjacent to $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$. We denote by ($\mathcal{I}^{\leq k}_{(a)}$, ($\mathcal{I}^{>k}_{(a)}$ the union of $\mathcal{I}^{\leq k}$, $\mathcal{I}^{>k}$ with the required adjacent indices and compute the right-hand side tensors in (58) by
$G_{Y}\left(\mathcal{I}^{\leq k-1},:,\mathcal{I}^{>k},t\right)=G\left(Y\left(\mathcal{I}^{\leq k-1}_{(a)},:,\mathcal{I}^{>k}_{(a)},t\right),t\right),\qquad k=1,2,\ldots,d.$ (59)
Computing (59) is efficient for any nonlinear $G$ that we can evaluate entry-wise, regardless of the low-rank structure in $G$. The evolution equations (58) have a clear computational advantage over the evolution equations of orthogonal dynamical low-rank approximation or step-truncation methods [42], which require a low-rank representation of $G$ to be practical.
5.1.1 Constructing the low-rank solution in TT format
The (time-dependent) tensor cross interpolation (54) formula
$Y(i_{1},\ldots,i_{d},t)=\prod_{k=1}^{d-1}Y\left(\mathcal{I}^{\leq k-1}(t),i_{k},\mathcal{I}^{>k}(t),t\right)\left[Y\left(\mathcal{I}^{\leq k}(t),\mathcal{I}^{>k}(t),t\right)\right]^{-1}Y\left(\mathcal{I}^{\leq d-1}(t),i_{d},t\right),$ (60)
allows one to construct the solution to (57) at any time $t$ from the solution to the equations (58). However, using (60) directly can lead to numerical instability as the time-dependent $r_{k}\times r_{k}$ matrices $Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k},t\right)$ can be ill-conditioned. Hereafter we describe a more robust method for computing $Y(t)$ by orthogonalization, omitting the dependence of tensors on $t$ to simplify notation. Take QR-decompositions
$\left[\bm{Y}\left(\mathcal{I}^{\leq k-1},:,\mathcal{I}^{>k}\right)\right]^{\langle l\rangle}=\bm{Q}_{k}^{\langle l\rangle}\bm{R}_{k},\qquad k=1,2,\ldots,d-1,$ (61)
to write
$Y\left(\mathcal{I}^{\leq k-1},i_{k},\mathcal{I}^{>k}\right)=Q_{k}(i_{k})\bm{R}_{k},\qquad Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k}\right)=\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:\right)\bm{R}_{k},$ (62)
where $\bm{l}_{\leq k}$ is defined in (45). Then substitute (62) into (60) to obtain
$\begin{split}Y(i_{1},\ldots,i_{d})&=\prod_{k=1}^{d-1}Q_{k}(i_{k})\bm{R}_{k}\left[\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:\right)\bm{R}_{k}\right]^{-1}Y\left(\mathcal{I}^{\leq d-1},i_{d}\right)\\ &=\prod_{k=1}^{d-1}Q_{k}(i_{k})\left[\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:\right)\right]^{-1}Y\left(\mathcal{I}^{\leq d-1},i_{d}\right),\end{split}$ (63)
Computing $Y$ via (63) instead of (60) yields a more stable numerical algorithm as the matrices $\widehat{\bm{Q}}_{k}=\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:\right)$ are related to the orthogonal bases $\bm{U}_{\leq k}$ from which the multi-index sets $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ were obtained with the TT-cross-DEIM index selection algorithm and therefore are have smaller condition number than $Y\left(\mathcal{I}^{\leq k},\mathcal{I}^{>k}\right)$. The improvement in condition number is indeed verified by our numerical experiments as shown in Figure 5(d).
5.1.2 Discrete-time TT-cross integrator
Let us describe one step of the TT-cross time integration scheme from time $t_{0}$ to $t_{1}=t_{0}+\Delta t$ starting from the rank-$\bm{r}$ TT representation
$Y(t_{0})=C_{1}(t_{0})C_{2}(t_{0})\cdots C_{d}(t_{0}).$ (64)
First compute the indices $\left\{\mathcal{I}^{\leq k}(t_{0}),\mathcal{I}^{>k}(t_{0})\right\}$ for the interpolatory projector defining the dynamical low-rank evolution equation (57) using the TT-cross-DEIM algorithm. Then integrate the evolution equations (58) from time $t_{0}$ to $t_{1}$ using an explicit time stepping scheme with multi-indices fixed at time $t_{0}$, e.g., Euler forward yields
$\begin{split}Y\left(\mathcal{I}^{\leq k-1}(t_{0}),:,\mathcal{I}^{>k}(t_{0}),t_{1}\right)=Y\left(\mathcal{I}^{\leq k-1}(t_{0}),:,\mathcal{I}^{>k}(t_{0}),t_{0}\right)+\Delta tG_{Y}\left(\mathcal{I}^{\leq k-1}(t_{0}),:,\mathcal{I}^{>k}(t_{0}),t_{0}\right),\end{split}$ (65)
for all $k=1,2,\ldots,d$. Use the result of explicit time integration (65) to construct TT-cores for the solution at time $t_{1}$
$Y(t_{1})=C_{1}(t_{1})C_{2}(t_{1})\cdots C_{d}(t_{1}),$ (66)
with the QR-stabilized procedure described in (61)-(63), i.e.,
$\begin{split}&C_{k}(i_{k},t_{1})=Q_{k}(i_{k},t_{1})\left[\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k}(t_{0}),:,t_{1}\right)\right]^{-1},\quad k=1,2,\ldots,d-1,\\ &C_{d}(t_{1})=Y\left(\mathcal{I}^{\leq d-1}(t_{0}),:,t_{1}\right).\end{split}$ (67)
This completes one step of the TT-cross time integration scheme.
Computational cost
To simplify the operation count of one step of the TT-cross integrator, we assume that $r_{k}=r$ and $n_{k}=n$ for all $k=1,2,\ldots,d$. As shown in Section 4, the TT-cross-DEIM algorithm used to obtain the indices $\{\mathcal{I}^{\leq k},\mathcal{I}^{>k}\}$ requires $\mathcal{O}(dnr^{3})$ operations. Preparing the tensors $Y(\mathcal{I}^{\leq k-1},:\mathcal{I}^{>k})$ required to take the explicit time step (65) requires $d-1$ matrix multiplications with matrices of size $r\times r$ and $r\times(nr)$ for a number of operations scaling as $\mathcal{O}(dnr^{3})$. For many $G$ that can be evaluated entry-wise (e.g., entry-wise nonlinearities), the cost of evaluating $G_{Y}(^{\leq k-1},:,\mathcal{I}^{>k};t)$ is on the same order of computing the subtensors $Y(\mathcal{I}^{\leq k-1},:\mathcal{I}^{>k})$, i.e., $\mathcal{O}(dnr^{3})$. Finally reconstructing the tensor cores at time $t_{1}$ in (67) requires $d-1$ QR-decompositions of matrices with size $r\times(nr)$ and inverting $d-1$ matrices of size $r\times r$ with total cost scaling as $\mathcal{O}(dnr^{3})$. Thus the computational cost of one step of the TT-cross time integrator scales as $\mathcal{O}(dnr^{3})$. We note that it is not strictly necessary to perform these steps at every time step as multi-indices can be reused over many time steps, provided the condition of $\bm{Q}_{k}^{\langle l\rangle}\left(\bm{l}_{\leq k},:,t\right)$ remains under control. If the computation of new indices is not required then (65) can be iterated for a number of time steps at a cost of $\mathcal{O}(dnr^{2})$ operations before computing new indices.
5.2 Interpolatory projector-splitting integrator
Next, we directly applying a splitting integrator to the dynamical low-rank evolution equation (57) to obtain a generalization of the orthogonal projector-splitting integrator introduced in [32] for TTs. Just like the orthogonal projector-splitting integrator, the oblique projector-splitting integrator is robust to small singular values and one step of the integrator can be implemented as an efficient sweeping algorithm over the TT-cores of the solution. Using the expression (14) for the oblique tangent space projector the right-hand side of the dynamical low-rank evolution equation (57) can be written as a sum of $12d-$ 1 terms
$P_{Y}G(Y,t)=\sum_{j=1}^{d-1}P_{j}^{+}G(Y,t)-P_{j}^{-}G(Y,t)+P_{d}^{+}G(Y,t),$ (68)
where $P_{j}^{+}=P_{\leq j-1}P_{>j}$ and $P_{j}^{-}=P_{\leq j}P_{>j}$. Integrating (57) from time $t_{0}$ to $t_{1}=t_{0}+\Delta t$ with first order Lie-Trotter splitting requires solving the $12d-$ 1 substeps
$\frac{dY_{1}^{+}(t)}{dt}$ $=P_{1}^{+}G\left(Y_{1}^{+},t\right),$ $Y_{1}^{+}(t_{0})=Y(t_{0}),$ (69)
$\frac{dY_{1}^{-}(t)}{dt}$ $=-P_{1}^{-}G\left(Y_{1}^{-},t\right),\quad$ $Y_{1}^{-}(t_{0})=Y_{1}^{+}(t_{1}),$
[…]
$\frac{dY_{j}^{+}(t)}{dt}$ $=P_{j}^{+}G\left(Y_{j}^{+},t\right),$ $Y_{j}^{+}(t_{0})=Y_{j-1}^{-}(t_{1}),$
$\frac{dY_{j}^{-}(t)}{dt}$ $=-P_{j}^{-}G\left(Y_{j}^{-},t\right),$ $Y_{j}^{-}(t_{0})=Y_{j}^{+}(t_{1}),$
[…]
$\frac{dY_{d}^{+}(t)}{dt}$ $=P_{d}^{+}G\left(Y_{d}^{+},t\right),$ $Y_{d}^{+}(t_{0})=Y_{d-1}^{-}(t_{1}),$
in consecutive order to obtain the approximate solution $Y(t_{1})=Y_{d}^{+}(t_{1})$ at time $t_{1}$. Note that the projectors $P_{j}^{+},P_{j}^{-}$ depend on solutions to each substep $Y_{j}^{+}(t)$ or $Y_{j}^{-}(t)$ and thus are time-dependent. In the case of orthogonal tangent space projector-splitting it was shown in [32, Theorem 4.1] that $P_{j}^{+},P_{j}^{-}$ can be kept constant during each substep and each of the differential equations (69) can be solved exactly by updating a single TT-core. The following Theorem is an analogous result for the oblique projector-splitting integrator. We suppress the dependence of the multi-indices $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ on $t$ although it is assumed that such indices are selected at each time $t$ so that the interpolatory projectors (14) (and therefore the oblique tangent space projector (16)) are well-conditioned.
Theorem 5.1
Each split differential equation in (69) is solved exactly with time-independent projectors $P_{j}^{+}$ and $P_{j}^{-}$ at $Y_{j}^{+}(t_{0})$ and $Y_{j}^{-}(t_{0})$, respectively. Moreover, if $Y_{j}^{+}(t_{0})$ has the TT representation
$Y_{j}^{+}(t_{0})=U_{\leq j-1}\left[U_{j}\bm{S}_{j}\right]V_{>j}$ (70)
then
$Y_{j}^{+}(t)=U_{\leq j-1}K_{j}(t)V_{>j},$
where
$\frac{dK_{j}(t)}{dt}=\left[U_{\leq j-1}\left(\mathcal{I}^{\leq j-1},:\right)\right]^{-1}G_{j}^{+}\left(\mathcal{I}^{\leq j-1},:,\mathcal{I}^{>j},t\right)\left[V_{>j}\left(:,\mathcal{I}^{>j}\right)\right]^{-\text{T}},\quad K_{j}(t_{0})=U_{j}\bm{S}_{j},$ (71)
and $G_{j}^{+}(t)=G\left(Y_{j}^{+}(t),t\right)$.
Similarly if $Y_{j}^{-}(t_{0})$ has the TT representation $Y_{j}^{-}(t_{0})=U_{\leq j}\bm{S}_{j}(t_{0})V_{>j}$ then
$Y_{j}^{-}(t)=U_{\leq j}\bm{S}_{j}(t)V_{>j},$
where
$\frac{d\bm{S}_{j}(t)}{dt}=-\left[U_{\leq j}\left(\mathcal{I}^{\leq j},:\right)\right]^{-1}G_{j}^{-}\left(\mathcal{I}^{\leq j},\mathcal{I}^{>j},t\right)\left[V_{>j}\left(:,\mathcal{I}^{>j}\right)\right]^{-\text{T}},$ (72)
and $G_{j}^{-}=G\left(Y_{j}^{-}(t),t\right)$.
Proof: The proof follows the same approach as the corresponding result for the orthogonal projector-splitting integrator. First recall that we have shown in the proof of Proposition 3.1 that $P_{\leq j-1}P_{>j}$ maps onto a tangent space of $\mathcal{M}_{\bm{r}}$ at each time $t$. This ensures that $Y_{j}^{+}(t)$ belongs to $\mathcal{M}_{\bm{r}}$ for all $t$ and therefore admits a time-dependent orthogonalized rank-$\bm{r}$ TT decomposition of the form
$Y_{j}^{+}(t)=U_{\leq j-1}(t)K_{j}(t)V_{>j}(t).$ (73)
Substituting (73) into (69) and using the product rule we obtain
$\begin{split}&\frac{dU_{\leq j-1}(t)}{dt}K_{j}(t)V_{>j}(t)+U_{\leq j-1}(t)\frac{dK_{j}(t)}{dt}V_{>j}(t)+U_{\leq j-1}(t)K_{j}(t)\frac{dV_{>j}(t)}{dt}\\ &\quad=P_{\leq j-1}P_{>j}G\left(Y_{j}^{+}(t),t\right)\\ &\quad=U_{\leq j-1}(t)\delta C_{j}(t)V_{>j}(t),\end{split}$ (74)
where we used (21) to obtain the third line with
$\begin{split}\delta C_{j}(t)=\left[U_{\leq j-1}\left(\mathcal{I}^{\leq j-1},:,t\right)\right]^{-1}G_{j}^{+}\left(\mathcal{I}^{\leq j-1},:,\mathcal{I}^{>j},t\right)\left[V_{>j}\left(:,\mathcal{I}^{>j},t\right)\right]^{-\text{T}}\end{split}$ (75)
Equation (74) is solved exactly by setting $dU_{\leq j-1}(t)/dt=0$, $dV_{>j}(t)/dt=0$ and $dK_{j}(t)/dt=\delta C_{j}(t)$ and from the initial condition $Y_{j}^{+}(t_{0})$ in (70) we obtain
$\begin{split}&U_{\leq j-1}(t)=U_{\leq j-1},\quad V_{>j}(t)=V_{>j},\quad K_{j}(t_{0})=U_{j}\bm{S}_{j},\end{split}$ (76)
proving the result for $Y_{j}^{+}(t)$. The proof of the assertion for $Y_{j}^{-}(t)$ is similar. $\square$
Similar to the TT-cross evolution equations (58), computing the right-hand side of the differential equation (71) requires evaluating $G_{j}^{+}$ at a subset of $r_{j-1}n_{j}r_{j}$ indices and computing the right-hand side of (72) requires evaluating $G_{j}^{-}$ at a subset of $r_{j}^{2}$ indices. These evaluations are efficient for any $G$ that can be evaluated entry-wise and do not require $G$ to have any low-rank structure. The differential equations (71) and (72) also feature inverses of $r_{j}\times r_{j}$ matrices $U_{\leq j}\left(\mathcal{I}^{\leq j},:\right)$ and $V_{>j}\left(:,\mathcal{I}^{>j}\right)$. These matrices define the interpolatory projectors (14) and we select the multi-indices $\mathcal{I}^{\leq j},\mathcal{I}^{>j}$ at each time $t$, e.g., with the TT-cross-DEIM, to ensure that their condition number is small.
5.2.1 Sweeping algorithm for interpolatory projector-splitting integrator
One complete step of the splitting integrator can be implemented as an efficient sweeping algorithm that updates the TT-cores from left to right and computes the necessary index sets for oblique projection. To describe one step of the sweeping algorithm from time $t_{0}$ to $t_{1}=t_{0}+\Delta t$, suppose we have an orthogonalized TT decomposition of solution at time $t_{0}$ of the form
$Y(t_{0})=U_{1}(t_{0})S_{1}(t_{0})V_{>1}(t_{0}),$ (77)
and the multi-indices $\mathcal{I}^{>j}(t_{0})$ defining interpolatory projectors onto the bases $V_{>j}(t_{0})$ for $j=1,2,\ldots,d-1$. The sweeping algorithm solves the equations in (69) sequentially by updating the solution TT-cores $U_{j}(t_{0})$ to $U_{j}(t_{1})$ and then computes the indices $\mathcal{I}^{\leq j}(t_{1})$ for the oblique projectors (14) onto the updated bases $\bm{U}_{\leq j}(t_{1})$ required for the next step in the sweep.
To begin we apply Theorem 5.1 to solve the first differential equation in (69) by integrating
$\frac{dK_{1}(t)}{dt}=G_{1}^{+}\left(:,\mathcal{I}^{>1}(t_{0}),t\right)\left[V_{>1}\left(:,\mathcal{I}^{>1}(t_{0}),t_{0}\right)\right]^{-\text{T}},\qquad K_{1}(t_{0})=U_{1}(t_{0})\bm{S}_{1}(t_{0}),$ (78)
from $t_{0}$ to $t_{1}$. The solution $Y_{1}^{+}(t_{1})=K_{1}(t_{1})V_{>1}(t_{0})$ is the starting value $Y_{1}^{-}(t_{0})=Y_{1}^{+}(t_{1})$ for the second equation in (69). We then prepare $Y_{1}^{-}(t_{0})$ for the application of Theorem 5.1 by decomposing $K_{1}(t_{1})=U_{1}(t_{1})\bm{R}_{1}(t_{1})$ to obtain $Y_{1}^{-}(t_{0})=U_{1}(t_{1})\bm{R}_{1}(t_{1})V_{>1}(t_{0})$ where $U_{1}(t_{1})$ is left-orthogonal (see Section 2.1) and compute indices $\mathcal{I}^{\leq 1}(t_{1})=\texttt{DEIM}(\bm{U}_{\leq 1}(t_{1}))$. Now we can apply Theorem 5.1 to solve the second differential equation in (69) by integrating
$\frac{d\bm{S}_{1}(t)}{dt}=-\left[U_{\leq 1}\left(\mathcal{I}^{\leq 1}(t_{1}),:,t_{1}\right)\right]^{-1}G_{1}^{-}\left(\mathcal{I}^{\leq 1}(t_{1}),\mathcal{I}^{>1}(t_{0}),t\right)\left[V_{>1}\left(:,\mathcal{I}^{>1}(t_{0}),t_{0}\right)\right]^{-\text{T}},\quad\bm{S}_{1}(t_{0})=\bm{R}_{1}(t_{1}),$ (79)
from time $t_{0}$ to $t_{1}$ to obtain the solution $Y_{1}^{-}(t_{1})=U_{1}(t_{1})\bm{S}_{1}(t_{1})V_{>1}(t_{0})$. The algorithm proceeds recursively with step $j$ of the sweep described below.
Computing $Y_{j}^{+}(t_{1})$
The starting value $Y_{j}^{+}(t_{0})=Y_{j-1}^{-}(t_{1})$ is available in the form
$Y_{j}^{+}(t_{0})=U_{\leq j-1}(t_{1})\bm{S}_{j-1}(t_{1})V_{>j-1}(t_{0}),$ (80)
from the computation of $Y_{j-1}^{-}(t_{1})$, as are the multi-index sets $\mathcal{I}^{\leq j-1}(t_{1})$ and $\mathcal{I}^{>j}(t_{0})$. To apply Theorem 5.1 we write (80) as
$Y_{j}^{+}(t_{0})=U_{\leq j-1}(t_{1})\left[\bm{S}_{j-1}(t_{1})V_{j}(t_{0})\right]V_{>j}(t_{0}),$ (81)
and then integrate
$\begin{split}\frac{dK_{j}(t)}{dt}=\left[U_{\leq j-1}\left(\mathcal{I}^{\leq j-1}(t_{1}),:,t_{1}\right)\right]^{-1}G_{j}^{+}\left(\mathcal{I}^{\leq j-1}(t_{1}),:,\mathcal{I}^{>j}(t_{0}),t\right)\left[V_{>j}\left(:,\mathcal{I}^{>j}(t_{0}),t_{0}\right)\right]^{-\text{T}},\\ \qquad K_{j}(t_{0})=\bm{S}_{j-1}(t_{1})V_{j}(t_{0}),\end{split}$ (82)
from $t_{0}$ to $t_{1}$ to obtain the solution $Y_{j}^{+}(t_{1})=U_{\leq j-1}(t_{1})K_{j}(t_{1})V_{>j}(t_{0})$.
Computing $Y_{j}^{-}(t_{1})$
The starting value $Y_{j}^{-}(t_{0})=Y_{j}^{+}(t_{1})$ is available in the form
$Y_{j}^{-}(t_{0})=U_{\leq j-1}(t_{1})K_{j}(t_{1})V_{>j}(t_{0}),$ (83)
from the computation of $Y_{j}^{+}(t_{1})$ as are the multi-index sets $\mathcal{I}^{\leq j-1}(t_{1})$ and $\mathcal{I}^{>j}(t_{0})$. We prepare $Y_{j}^{-}(t_{0})$ for the application of Theorem 5.1 by decomposing $K_{j}(t_{1})=U_{j}(t_{1})\bm{R}_{j}(t_{1})$ where $U_{j}(t_{1})$ is left-orthogonal (see Section 2.1) which allows us to write the starting value (83) as
$\begin{split}Y_{j}^{-}(t_{0})=U_{\leq j}(t_{1})\bm{R}_{j}(t_{1})V_{>j}(t_{0}).\end{split}$ (84)
Then we obtain the multi-indices $\mathcal{I}^{\leq j}(t_{1})$ from $\mathcal{I}^{\leq j-1}(t_{1})$ and the TT-cores $U_{\leq j}(t_{1})$ with a substep of the TT-cross-DEIM algorithm as described in (44)-(46). Then by Theorem 5.1, integrating
$\frac{d\bm{S}_{j}(t)}{dt}=-\left[U_{\leq j}\left(\mathcal{I}^{\leq j}(t_{1}),:,t_{1}\right)\right]^{-1}G_{j}^{-}\left(\mathcal{I}^{\leq j}(t_{1}),\mathcal{I}^{>j}(t_{0}),t\right)\left[V_{>j}\left(:,\mathcal{I}^{>j}(t_{0}),t_{0}\right)\right]^{-\text{T}},\quad\bm{S}_{j}(t_{0})=\bm{R}_{j}(t_{1}),$ (85)
from time $t_{0}$ to $t_{1}$ yields the solution $Y_{j}^{-}(t_{1})=U_{\leq j}(t_{1})S_{j}(t_{1})V_{>j}(t_{0})$.
Iterating these steps until we obtain $Y_{d}^{+}(t_{1})=U_{\leq d-1}(t_{1})K_{d}(t_{1})=Y(t_{1})$ completes one step of the first-order splitting integrator. To take another time step the TT representation of $Y(t_{1})$ must be orthogonalized from right to left, i.e., (6) with $k=1$. During this orthogonalization procedure, the indices $\mathcal{I}^{>j}(t_{1})$ can be computed with the right-to-left TT-cross-DEIM sweep as described in Section 4.1. Similar to the orthogonal projector-splitting integrator, obtaining the second-order Strang projector-splitting integrator is straightforward by composing the Lie-Trotter integrator with its adjoint. In this case the forward sweep described above is performed with step-size $\Delta t/2$ and is then followed by a backward sweep also with step-size $\Delta t/2$. The oblique projector-splitting integrator has the same computational complexity as the TT-cross integrator and in fact the two methods are closely related. The TT-cross method integrates TT-cores and then computes an oblique projection. Meanwhile the projector-splitting method computes the oblique projection first and then integrates TT-cores.
5.3 Rank-adaptive time integration
The solution to (2) is often not accurately represented on a tensor manifold $\mathcal{M}_{\bm{r}}$ with constant rank for all $t\in[0,T]$. Therefore the dynamical low-rank integrators must be able to decrease or increase the solution rank during time integration. To decrease the solution rank we use the TT-SVD truncation algorithm at each time $t$ which requires $d-1$ orthogonal representations (6) of the solution. Such orthogonalizations are required for the TT-cross-DEIM index selection algorithm and thus rank decrease can be performed during time integration with either the TT-cross or interpolatory projector-splitting algorithms at no additional computational cost.
To increase the $k$th component of the TT solution rank during integration with the TT-cross integrator we modify Algorithm 2 to sample $\hat{r}_{k}>r_{k}$ indices $\bm{l}_{\leq k}$ from the left singular vectors (45) and $\hat{r}_{k}$ indices $\bm{l}_{>k}$ from right singular vectors in (49) by augmenting the DEIM indices with additional indices selected by another sparse index selection algorithm, e.g., GappyPOD+E [39]. From the $\bm{l}_{\leq k},\bm{l}_{>k}$ we construct $\mathcal{I}^{\leq k},\mathcal{I}^{>k}$ in (46),(50) each with $\hat{r}_{k}$ indices. We then integrate the solution $Y(t)$ forward in time on the manifold $\mathcal{M}_{\hat{\bm{r}}}$ using the equations (58). It is well-known that the solution $Y(t)$ with rank $\bm{r}$ belongs to the boundary of the higher rank manifold $\mathcal{M}_{\hat{\bm{r}}}$ where the tangent space is not well-defined [51]. Nevertheless, the evolution equations (58), which define the interpolatory tangent space projection, are well-defined on the boundary of $\mathcal{M}_{\hat{\bm{r}}}$. These equations allow us to integrate $Y(t)$ forward in time on $\mathcal{M}_{\hat{\bm{r}}}$ thereby increasing the solution rank. To increase the $k$th component of the TT solution rank during integration with the projector-splitting integrator we add new (orthogonal) basis vectors to the TT cores with zero singular value and then sample indices from this augmented basis and apply the projector-splitting integrator to the augmented solution. Once again adding new basis vectors with zero singular values places the approximate solution on the boundary of a higher rank manifol $\mathcal{M}_{\hat{\bm{r}}}$. However the projector-splitting integrator is robust to zero singular values and allows us to integrate the solution off of the boundary of the low-rank manifold.
A simple criterion for determining when to increase the $k$th component of the TT-rank vector is based on the singular values $\{\sigma_{k}(\alpha_{k},t)\}_{\alpha_{k}=1}^{r_{k}}$ of the unfolding matrix $\bm{Y}^{\langle k\rangle}$. We select the rank to ensure that the relative size of the smallest singular value
$\epsilon_{k}(t)=\frac{\sigma_{k}(r_{k},t)}{\sqrt{\displaystyle\sum_{\alpha_{k}=1}^{r_{k}}\sigma_{k}(\alpha_{k},t)^{2}}},\qquad k=1,2,\ldots,d-1$ (86)
remains in a desired range $\epsilon_{l}\leq\epsilon_{k}(t)\leq\epsilon_{u}$. This criterion is an adaptation of the rank-adaptive criterion proposed in [15] for matrix differential equations and subsequently generalized to the Tucker format [21], to the TT format.
6 Numerical examples
We now apply the proposed dynamical low-rank collocation methods to several tensor differential equations (2) arising from the discretization of partial differential equations (1) and compare the accuracy and efficiency with existing time integration schemes on tensor manifolds. We measure the accuracy of the low-rank approximations $Y(t)$ to the solution $X(t)$ of (2) in the relative Frobenius norm
$E(t)=\frac{\left\|Y(t)-X(t)\right\|_{F}}{\left\|X(t)\right\|_{F}}.$ (87)
To illustrate the key differences between the proposed interpolatory projector-splitting integrator the orthogonal projector-splitting integrator [31] we begin with a two-dimensional example.
6.1 2D Vlasov-Poisson equation
We consider the Vlasov-Poisson equation
$\begin{cases}\displaystyle\frac{\partial u(x,v,t)}{\partial t}+v\displaystyle\frac{\partial u(x,v,t)}{\partial x}+E(x)\displaystyle\frac{\partial u(x,v,t)}{\partial v}=0\\ u(x,v,0)=u_{0}(x,v),\end{cases}$ (88)
from [24, Example 4.4] with initial condition $f(x,v,0)=\exp(-20(x^{2}+v^{2}))$, electric field $E(x)=0.5\sin(\pi x)$ and $x\in\Omega_{x}=[-1,1],v\in\Omega_{v}=[-1,1]$. Discretizing $\Omega_{x}$ and $\Omega_{v}$ using $n=64$ points and approximating derivatives with a Fourier pseudo-spectral method [26] we obtain a semi-discrete version of the Vlasov-Poisson equation (88) in the form of a differential equation (2) with $d=2$, i.e., a matrix differential equation. To assess the accuracy of the low-rank tensor integrators, we computed a reference solution $\bm{X}(t)$ by integrating the full matrix differential equation to the semi-discrete equation with the four-stage explicit Runge-Kutta (RK4) method using time step-size $\Delta t=10^{-3}$.
We compared the TT-cross integrator presented in Section 5.1 with a step-truncation method using SVD-based truncation (ST-SVD). For both integrators we used Adams-Bashfort 2 with step-size $\Delta t=10^{-3}$. We utilized the rank-adaptive mechanism described in Section 5.3 for TT-cross with parameter $\epsilon_{l}=10^{-7}$. For the ST-SVD solution we set relative truncation tolerance $\delta=10^{-7}$ at each time step allowing the solution rank to adapt in time accordingly. In Figure 4(b) we plot the rank of the TT-cross and ST-SVD solutions versus time. The rank grows rapidly during time integration which allows us to assess the robustness of the rank-adaptive mechanism for the TT-cross integrator. In Figure 4 we plot the relative error of the TT-cross and ST-SVD solutions in the Frobenius norm versus time. We observe that the TT-cross solution is more accurate than the ST-SVD solution due to the TT-cross solution rank being slightly larger than the rank of the ST-SVD solution at each step. The error of the TT-cross solution remains controlled during time integration, demonstrating the effectiveness of the rank-adaptive mechanism.
Next we compared the interpolatory projector-splitting integrator (i-PS) presented in Section 5.2 with the orthogonal projector-splitting integrator (o-PS) introduced in [31]. In this case ($d=2$), the sweeping algorithm (see Section 5.2.1) for one step of the projector-splitting integrator has three substeps commonly referred to as $K$-, $S$-, and $L$-step. Beginning from the rank-$r$ decomposition of the approximate solution $\bm{Y}(t_{0})=\bm{U}(t_{0})\bm{S}(t_{0})\bm{V}^{\text{T}}(t_{0})$ at time $t_{0}$, these three steps yield low-rank factors for the approximate solution $\bm{Y}(t_{1})=\bm{U}(t_{1})\bm{S}(t_{1})\bm{V}^{\text{T}}(t_{1})$ at time $t_{1}=t_{0}+\Delta t$. The substeps for interpolatory projector-splitting integrator are as follows.
  • 1.
    K-step: update $\bm{U}(t_{0})\to\bm{U}(t_{1})$ and $\bm{S}(t_{0})\to\bm{R}(t_{1})$.
    Compute interpolation indices $\mathcal{J}=\texttt{DEIM}(\bm{V}(t_{0}))$. Then integrate the $n_{1}\times r$ differential equation
    $\frac{d\bm{K}(t)}{dt}=\bm{G}_{\bm{K}}\left(:,\mathcal{J},t\right)\left[\bm{V}\left(\mathcal{J},:,t_{0}\right)\right]^{-\text{T}},\qquad\bm{K}(t_{0})=\bm{U}(t_{0})\bm{S}(t_{0}),$ (89)
    where $\bm{G}_{\bm{K}}(t)=G\left(\bm{K}(t)\bm{V}(t_{0})^{\text{T}},t\right)$ from $t_{0}$ to $t_{1}$, and perform a QR-decomposition $\bm{K}(t_{1})=\bm{U}(t_{1})\bm{R}(t_{1})$.
  • 2.
    S-step: update $\bm{R}(t_{1})\to\tilde{\bm{S}}(t_{1})$.
    Compute interpolation indices $\mathcal{I}=\texttt{DEIM}(\bm{U}(t_{1}))$. Then integrate the $r\times r$ matrix differential equation
    $\frac{d\tilde{\bm{S}}(t)}{dt}=-\left[\bm{U}\left(\mathcal{I},:,t_{1}\right)\right]^{-1}\bm{G}_{\bm{S}}\left(\mathcal{I},\mathcal{J},t\right)\left[\bm{V}\left(\mathcal{J},:,t_{0}\right)\right]^{-\text{T}},\qquad\tilde{\bm{S}}(t_{0})=\bm{R}(t_{1}),$ (90)
    where $\bm{G}_{\bm{S}}(t)=G\left(\bm{U}(t_{1})\bm{S}(t)\bm{V}(t_{0})^{\text{T}},t\right)$ from $t_{0}$ to $t_{1}$.
  • 3.
    L-step: update $\bm{V}(t_{0})\to\bm{V}(t_{1})$ and $\tilde{\bm{S}}(t_{1})\to\bm{S}(t_{1})$.
    Integrate the $n_{2}\times r$ matrix differential equation
    $\frac{d\bm{L}(t)}{dt}=\bm{G}_{\bm{L}}\left(\mathcal{I},:,t\right)^{\text{T}}\left[\bm{U}\left(\mathcal{I},:,t_{1}\right)\right]^{-\text{T}},\qquad\bm{L}(t_{0})=\bm{V}(t_{0})\tilde{\bm{S}}(t_{1})^{\text{T}},$ (91)
    where $\bm{G}_{\bm{L}}(t)=G\left(\bm{U}(t_{1})\bm{L}(t)^{\text{T}},t\right)$ from time $t_{0}$ to $t_{1}$, and perform a QR-decomposition $\bm{L}(t_{1})=\bm{V}(t_{1})\bm{S}(t_{1})^{\text{T}}$.
                 (a)                                                                                   (b)
                       (c)                                                                                   (d)
Figure 4: Low-rank approximations to the solution of the two-dimensional Vlasov-Poisson equation (88). (a) Relative error versus time of TT-cross and ST-SVD solutions. The TT-cross solution was computed using rank-adaptive singular value threshold $\epsilon_{l}=10^{-7}$ and the ST-SVD solution was computed using truncation threshold $\delta=10^{-7}$. (b) Rank versus time of the rank-adaptive TT-cross and ST-SVD solutions. (c) Relative error versus time of solutions computed with interpolatory and orthogonal projector-splitting using rank-adaptive singular value threshold $\epsilon_{l}=10^{-7}$. (d) Rank versus time of the rank-adaptive solutions computed with interpolatory and orthogonal projector splitting integrators.
The interpolatory projector-splitting integrator requires evaluating the tensors $\bm{G}_{\bullet}$ at only a subset of $nr$ indices in the $K$- and $L$-step and $r^{2}$ indices in the $S$-step. These evaluations can be performed efficiently for any $G$ that can be evaluated entry-wise. Meanwhile the corresponding steps of the orthogonal projector-splitting integrator (summarized in [6]) involve inner products between the tensors $\bm{G}_{\bullet}$ that only efficient to compute when $G$ has low-rank structure.
For the interpolatory and orthogonal projector-splitting integrators we used step-size $\Delta t=10^{-3}$ and solved the differential equations in the $K$, $S$, and $L$-steps with RK4. We also used the rank-adaptive mechanism described in Section 5.3 with parameter $\epsilon_{l}=10^{-7}$ for both solutions. In Figure 4(b) we plot the solution ranks versus time. Both solutions have the same rank until approximately $t=0.7$ when the i-PS solution rank becomes slightly smaller than the o-PS solution rank. In Figure 4(a) we plot the relative errors in the Frobenius norm versus time. The error of the i-PS and o-PS solutions is similar until around $t=0.4$, at which point the i-PS solution becomes slightly less accurate. This difference in accuracy is due to the i-PS method computing a quasi-optimal projection onto the tangent space, while the o-PS method computes the optimal projection at each time step. In addition, the slight difference in rank of the solutions also contributes to the difference in accuracy.
6.2 3D Allen-Cahn equation
The Allen-Cahn equation is a reaction-diffusion PDE that models phase separation in multi-component alloy systems [1, 28]. A simple form of such equation features a Laplacian and a cubic non-linearity
$\begin{cases}\displaystyle\frac{\partial u(\bm{x},t)}{\partial t}=\alpha\Delta u(\bm{x},t)+u(\bm{x},t)-u(\bm{x},t)^{3},\\ u(\bm{x},0)=u_{0}(\bm{x}).\end{cases}$ (92)
We consider the spatial domain $\Omega=[0,2\pi]^{3}$ with periodic boundary conditions, initial condition $u_{0}(x_{1},x_{2},x_{3})=g(x_{1},x_{2},x_{3})-g(2x_{1},x_{2},x_{3})+g(x_{1},2x_{2},x_{3})-g(x_{1},x_{2},2x_{3})$ where
$g(x_{1},x_{2},x_{3})=\frac{\left(e^{-\tan(x_{1})^{2}}+e^{-\tan(x_{2})^{2}}+e^{-\tan(x_{3})^{2}}\right)\sin(x_{1}+x_{2}+x_{3})}{1+e^{|\csc(-x_{1}/2)|}+e^{|\csc(-x_{2}/2)|}+e^{|\csc(-x_{3}/2)|}},$ (93)
and diffusion parameter $\alpha=0.1$. Discretizing $\Omega$ using $n=64$ points in each dimension and approximating derivatives with a Fourier pseudo-spectral method [26], we obtain a semi-discrete version of the Allen-Cahn equation in the form of (2). To assess the accuracy of the low-rank tensor integrators, we computed a reference solution $X(t)$ by integrating the full tensor solution ($64^{3}$ degrees of freedom) to the semi-discrete equation with the four-stage explicit Runge-Kutta (RK4) method using time step-size $\Delta t=10^{-3}$.
                 (a)                                                                                      (b)
                       (c)                                                                                        (d)
Figure 5: Low-rank approximations to the solution of the three-dimensional Allen-Cahn equation (92) computed with the TT-cross and ST-SVD methods. The ranks were determined using different truncation tolerances $\delta=10^{-4},10^{-6},10^{-10}$ in the ST-SVD method. (a) Relative error in the Frobenius norm versus time. (b) 1-norm of the TT-rank vector versus time. (c) Relative error of interpolatory (i-proj) and orthogonal (o-proj) projections onto the tensor manifold tangent space versus time. (d) Condition number of the matrices non-orthogonalized matrices in (60) and the corresponding orthogonalized matrices in (63) used to construct the TT-cross solution at each time step.
We compared the TT-cross integrator presented in Section 5.1 with the step-truncation SVD (ST-SVD) integrator [42] using three different relative truncation tolerances $\delta=10^{-4},10^{-6},10^{-10}$ for determining the solution rank at each time step. We set the solution rank in the TT-cross simulations equal to the ranks obtained from the ST-SVD simulations with truncation tolerances in order to compare the methods for solutions of the same rank. The rank decrease was performed using TT-SVD truncation and the rank increase by sampling more tensor cross indices than singular vectors using the GappyPOD+E algorithm [39] as described in Section 5.3. Time integration for both ST-SVD and TT-cross was performed with Adams-Bashforth 2 and step-size $\Delta t=10^{-3}$.
In Figure 5(b), we plot the 1-norm of the ST-SVD and TT-cross solution ranks. The smoothing effects due to diffusion in the Allen-Cahn equation cause the TT ranks to decay relatively quickly from time $t=0$ to time $t\approx 1.5$. In Figure 5(a), we plot the relative error measured in the Frobenius norm of the ST-SVD and TT-cross solutions versus time. The ST-SVD solution is more accurate than the TT-cross solution computed with the same rank, which is expected. Indeed, the ST-SVD method computes the best rank-$\bm{r}$ projection of the solution onto the low-rank manifold $\mathcal{M}_{\bm{r}}$ at each time step while the TT-cross method computes a quasi-optimal projection onto the tangent space of the manifold at each time step. When the rank of the TT solutions is large enough (in this case corresponding to $\delta=10^{-10}$), the time integration error dominates the low-rank approximation error and the ST-SVD and TT-cross methods produce solutions with the same accuracy. When the low-rank error dominates the time integration error ($\delta=10^{-4},10^{-6}$) we observe in Figure 5(a) that the ST-SVD is about half an order of magnitude more accurate than the TT-cross solution of the same rank for all ranks and at each time $t$. In Figure 5(c), we compare the accuracy of the interpolatory projection (i-proj) onto the tangent space computed from the TT-cross solution and the orthogonal projection (o-proj) onto the tangent space computed from the ST-SVD simulation. The orthogonal projection is more accurate than the interpolatory projection by approximately one order of magnitude or less at each time $t$. Similar to the difference in error between the solutions, the difference in error between the i-proj and o-proj is constant over all ranks and for all time $t$.
                     (a)                                                                                      (b)
Figure 6: Low-rank approximations to the solution of the three-dimensional Allen-Cahn equation (92) computed with the interpolatory and orthogonal projector-splitting integrators. Solutions are truncated at each time $t$ using tolerances TT-SVD with relative tolerance $\delta=10^{-6},10^{-10}$. (a) Relative error in the Frobenius norm versus time. (b) 1-norm of the TT-rank vector versus time.
The improved accuracy of the ST-SVD method over the TT-cross method comes at a significant computational cost due to the cubic nonlinearity in the Allen-Cahn equation (92). The reason is that the ST-SVD method requires computing a TT representation of $G(Y(t),t)$ at each time $t$, which is costly. Indeed, recall that standard algorithms for multiplying two TTs $Y_{1}$ and $Y_{2}$ with ranks $\bm{r}_{1}=\begin{bmatrix}r_{1}&\cdots&r_{1}\end{bmatrix}$ and $\bm{r}_{2}=\begin{bmatrix}r_{2}&\cdots&r_{2}\end{bmatrix}$ results in a TT $Y_{1}Y_{2}$ with rank equal to the Hadamard (element-wise) product of the two ranks $\bm{r}_{1}\circ\bm{r}_{2}$. These ranks are in general not optimal and to control the TT rank we perform a TT-SVD truncation requiring $\mathcal{O}(dn(r_{1}r_{2})^{3})$ operations. We used two TT-SVD truncations $\mathfrak{T}_{\delta}$ with relative accuracy $\delta$ to compute the cubic term
$(Y)^{3}=\mathfrak{T}_{\delta}^{\rm svd}\left(Y\ \mathfrak{T}_{\delta}^{\rm svd}\left(YY\right)\right),$ (94)
incurring a cost of $\mathcal{O}(dnr^{6})$ operations at each time $t$. It is possible to accelerate the computation $G(Y,t)$ by carrying out sums and products of TTs with approximate low-rank tensor arithmetic, black-box tensor cross approximation [14], or randomized algorithms [8]. However such algorithms introduce additional errors in the low-rank approximation that can be difficult to control. In comparison, the TT-cross integrator does not require $G(Y,t)$ in a low-rank form and instead evaluates $G(Y,t)$ at $\mathcal{O}(dnr^{2})$ indices. Thus the computational cost of the cubic nonlinearity for TT-cross is negligible compared to the $\mathcal{O}(dnr^{3})$ cost of the TT-cross-DEIM index selection algorithm and evaluating the subtensors of $Y(t)$ required to integrate the system of equations (58). The total CPU-time required to compute the ST-SVD solution with $\delta=10^{-4}$ was 3070 seconds, whereas the TT-cross method required 324 seconds. For $\delta=10^{-6}$, the CPU-time of ST-SVD amounted to 10670 seconds while the TT-cross method took 402 seconds.
We also compared the interpolatory projector-splitting (i-PS) integrator presented in Section 5.2 with the orthogonal projector-splitting (o-PS) integrator from [32] using two different truncation tolerances $\delta=10^{-6},10^{-10}$ on the singular values of the solutions. In both cases we used first-order Lie-Trotter splitting with time step-size $\Delta t=10^{-3}$ and solved each of the substeps in (69) with RK4. In Figure 6(a) we plot the error of the solutions computed with the i-PS and o-PS methods versus time. We observe that the i-PS method is less accurate than the o-PS method. This is expected since the i-PS method integrates $Y(t)$ on $\mathcal{M}_{\bm{r}}$ using a quasi-optimal tensor in the tangent space while the o-PS method integrates uses the optimal tensor in the tangent space. The difference in error is similar to the comparison of TT-cross and ST-SVD except for $t\in[0,5]$ in the simulations using $\delta=10^{-10}$ where the difference in error is significantly larger. In Figure 6(b) we plot the ranks of the i-PS and o-PS solutions versus time.
6.3 4D advection-diffusion-reaction equation
Finally we consider the advection-diffusion-reaction (ADR) equation
$\begin{cases}\displaystyle\frac{\partial u(\bm{x},t)}{\partial t}=\nabla\cdot\left(\bm{\mu}_{i}(\bm{x},t)u(\bm{x},t)\right)+\sigma\Delta u(\bm{x},t)+R(u)\\ u(\bm{x},0)=u_{0}(\bm{x}),\end{cases}$ (95)
where $R(u)$ is a nonlinear reaction term. We consider the spatial domain $\Omega=[0,2\pi]^{4}$ with periodic boundary conditions and set
$p_{0}(\bm{x})=\exp(\sin(x_{1})\sin(x_{2})\sin(x_{3})\sin(x_{4})),$ (96)
$R(u)=-0.1u/(1+u^{2})$, $\sigma=1/4$ and
$\bm{\mu}(\bm{x})=\frac{1}{2}\begin{bmatrix}g(x_{2},x_{3})\\ g(x_{3},x_{4})\\ g(x_{4},x_{1})\\ g(x_{2},x_{3})\end{bmatrix},$ (97)
where $g(x,y)=\exp(\sin(x)\cos(y))$. Discretizing $\Omega$ using $n=32$ points in each dimension and approximating derivatives with a Fourier pseudo-spectral method [26] we obtain a semi-discrete version of the ADR equation (95) in the form of (2). To assess the accuracy of the low-rank tensor integrators we computed a reference solution $X(t)$ by integrating the full tensor solution ($32^{4}$ degrees of freedom) to the semi-discrete equation with the four-stage explicit RK4 method using time step-size $\Delta t=10^{-3}$.
            $t=0$                                                                            $t=1$
Figure 7: The $(x_{1},x_{2})$-marginals of the reference solution to the four-dimensional ADR equation (95) at time $t=0$ and $t=1$.
We computed two approximate low-rank solutions on a TT manifold (8) with the step-truncation SVD method (ST-SVD) [42] using different relative truncation tolerances $\delta=10^{-6},10^{-8}$. Computing the ST-SVD solution requires a low-rank approximation of $G(Y(t),t)$ at each time $t$, which is challenging for the nonlinear ADR equation (95) as there are no reliable algorithms available for computing the fractional nonlinearity in the low-rank format. To compute the $G(Y(t),t)$, we construct the full tensor representation of the TT-SVD solution with $n^{4}$ degrees of freedom, compute the fractional nonlinearity, and then compress the result into a TT with a recursive SVD. This approach is of course not viable in higher dimensions but it allows us to compare our TT-cross solution with the ST-SVD method in this case which computes the best low-rank approximate solution at each time step.
The map $G$ obtained from discretizing (95) includes four coefficient tensors $c_{1},c_{2},c_{3},c_{4}\in\mathbb{R}^{n\times n\times n\times n}$ (resulting from the discretization of $g(x,y)$) that are not expressed in a low-rank format upon discretization of $G$. In order to compute $G(Y(t),t)$ in low-rank format at each time, we decomposed the four coefficient tensors in $G$ using TT-SVD compression with relative accuracy $\delta$. For $\delta=10^{-6}$ and $\delta=10^{-8}$ we obtained coefficient tensors of the same rank
$\begin{split}\text{TT-rank}(c_{1})&=\begin{bmatrix}1&1&12&1&1\end{bmatrix},\\ \text{TT-rank}(c_{2})&=\begin{bmatrix}1&1&1&12&1\end{bmatrix},\\ \text{TT-rank}(c_{3})&=\begin{bmatrix}1&13&13&13&1\end{bmatrix},\\ \text{TT-rank}(c_{4})&=\begin{bmatrix}1&1&12&1&1\end{bmatrix}.\\ \end{split}$ (98)
We computed $G(Y(t),t)$ in the ST-SVD method at each time step by taking products of the low-rank approximate coefficient tensors $c_{k}$ with the low-rank solution tensor $Y$ and then used TT-SVD truncation to compress the product. We then added the TT representation of the reaction term and applied TT-SVD truncation after adding two low-rank tensors in order to control tensor rank when computing $G(Y(t),t)$ at each time $t$. Time integration for the ST-SVD simulation was performed with AB2 and time step-size $\Delta t=10^{-3}$. In Figure 8(b) we plot the $1$-norm of the TT-rank of each ST-SVD solution versus time. We observe that the ranks of both solution increase until around $t=0.5$ and then stabilize for $t\in[0.5,1]$.
                    (a)                                                                                   (b)
Figure 8: Low-rank approximations to the solution of the four-dimensional ADR equation (95) computed with the TT-cross and ST-SVD integrators. The ranks were determined using different truncation tolerances $\delta=10^{-6},10^{-8}$ in the ST-SVD method. (a) Relative error in the Frobenius norm versus time. (b) 1-norm of the TT-rank vector versus time.
We then computed two approximate low-rank solutions on the TT manifold $\mathcal{M}_{\bm{r}}$ using the proposed TT-cross integrator. In order to compare the results with the ST-SVD simulations we set the solution ranks in the TT-cross simulations equal to the ranks obtained from the ST-SVD simulations with truncation tolerances $\delta=10^{-6},10^{-8}$. We computed the right-hand side of the TT-cross evolution equations (58) by simply evaluating the coefficient tensors at the indices determined by the TT-cross-DEIM Algorithm at each time step. The cost of computing the right hand-hand side for the TT-cross evolution equations is negligible compared to the $\mathcal{O}(dnr^{3})$ cost of the TT-cross-DEIM index selection algorithm and evaluating the subtensors of $Y(t)$ required to integrate the system of equations (58). In Figure 8(a) we compare the relative error in the Frobenius norm of the TT-cross solutions and the ST-SVD solutions. We observe that the TT-cross solutions are less accurate than the ST-SVD solutions of the same rank and the difference in accuracy is constant over all solution ranks and for all time $t$. This is expected as the ST-SVD method computes the best rank-$\bm{r}$ projection of the solution onto the TT manifold $\mathcal{M}_{\bm{r}}$ at each time step while the TT-cross method computes a quasi-optimal projection onto the tangent space of the low-rank manifold at each time step.
7 Conclusions
We introduced new general purpose dynamical low-rank methods for solving nonlinear differential equations on low-rank manifolds. The methods rely on a particular class of oblique projectors onto the tangent space with a low-rank manifold characterized by a cross-interpolation property. Such projectors collocate the differential equation on a low-rank tensor manifold and give rise to efficient time integration schemes that allow us to integrate differential equations defined by vector fields without low-rank structure on low-rank manifolds. To construct the oblique projections we introduced a new index selection algorithm based on the DEIM for constructing well-conditioned interpolatory projectors in the TT format. Furthermore, we showed that such indices also parameterize low-rank TT manifolds and their tangent spaces with cross interpolation. Our numerical results demonstrate that the oblique projections onto the tangent space yield good approximations on the low-rank manifold in the Frobenius norm that are efficiently computed for problems defined by vector fields without low-rank structure. Our proposed methods thus make dynamical low-rank approximation applicable to a broader class of differential equations and facilitate its use in various practical applications.
Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) program under the contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP-m1027.